### v1

# Tool-as-Interface: Learning Robot Policies from Human Tool Usage through Imitation Learning

Abstract
Tool use is critical for enabling robots to perform complex real-world tasks, and leveraging human tool-use data can be instrumental for teaching robots. However, existing data collection methods like teleoperation are slow, prone to control delays, and unsuitable for dynamic tasks. In contrast, human natural data—where humans directly perform tasks with tools—offers natural, unstructured interactions that are both efficient and easy to collect. Building on the insight that humans and robots can share the same tools, we propose a framework to transfer tool-use knowledge from human data to robots. Using two RGB cameras, our method generates 3D reconstruction, applies Gaussian splatting for novel view augmentation, employs segmentation models to extract embodiment-agnostic observations, and leverages task-space tool-action representations to train visuomotor policies. We validate our approach on diverse real-world tasks, including meatball scooping, pan flipping, wine bottle balancing, and other complex tasks. Our method achieves a 71% higher average success rate compared to diffusion policies trained with teleoperation data and reduces data collection time by 77%, with some tasks solvable only by our framework. Compared to hand-held gripper, UMI [16], our method cuts data collection time by 41%. Additionally, our method bridges the embodiment gap, improves robustness to variations in camera viewpoints and robot configurations, and generalizes effectively across objects and spatial setups.
Figure 1: Tool-as-Interface. We propose a scalable data collection and policy learning framework designed to transfer diverse, intuitive, and natural human data into effective visuomotor policies. The framework enables robots to learn robust policies that can operate effectively under challenging conditions, such as base and camera movement, and achieve high performance on a variety of complex manipulation tasks.
I Introduction
Tool use is essential to how humans interact with and transform their environment [67, 41]. For instance, humans use a pan to fry food and flip it, ensuring even cooking on both sides. Despite its significance, tool use beyond parallel jaw grippers remains underexplored in robotics, with research primarily focused on simpler tasks like grasping and pick-and-place operations [6, 46, 10, 36, 35]. In this paper, we focus on cost-effective data collection and efficient training of robot policies to rapidly acquire tool-use skills.
Imitation learning provides a promising pathway for robots to acquire tool-use skills by directly learning from human demonstrations [21, 26, 27, 29]. The paradigm excels in handling diverse tool-use tasks, as it bypasses the need for task-specific programming by relying solely on human demonstrations. However, its full potential hinges on addressing key challenges in collecting high-quality training data. Various teleoperation systems [33, 56, 8, 20, 29, 50, 84, 68, 39, 66, 45, 9] and hand-held grippers [65, 19, 53, 47, 43] have been developed to facilitate the collection of high-quality data. Teleoperation methods, such as kinematic replication and hand or body retargeting, show great potential [85, 22, 74, 62, 48]. However, their reliance on direct access to robot hardware limits both practicality and scalability. Hand-held grippers [57, 16] offer an alternative by enabling demonstrations in diverse environments. While they reduce dependency on robotic systems, their high costs and the technical expertise required for tasks like 3D printing and assembly restrict their accessibility to a specialized group of users.
To address these limitations, we turn to human interaction data — a natural, intuitive method through which humans interact with their environment during everyday activities without relying on external devices or specialized setups. Human interaction data refers to the natural process in which humans use their hands to operate tools and interact with and manipulate the environment freely. Unlike controlled demonstrations that require expensive hardware or meticulous preparation, human interaction data involves the spontaneous use of tools to interact with the environment. Human interaction data is an accessible, scalable, and cost-effective approach to data collection, requiring no prior knowledge or technical expertise, such as 3D printing or assembly, from participants. However, existing methods struggle to fully harness the potential of human interaction data. Key challenges include the embodiment gap and the reliance on single-view data, which limits the insights that can be drawn from interaction data [60, 4, 38, 5, 64, 42, 55].
Our framework addresses these challenges by leveraging the observation that humans and robots can share the same tools. We propose a novel approach that utilizes human interaction data to train robust and adaptable robot policies for diverse tool-use tasks (Figure 1). Our method minimizes reliance on expensive hardware, making data collection more scalable and accessible to non-experts. By capturing 3D information using two RGB cameras and generating 3D reconstructions, our method enables view-invariant policy learning through novel view augmentation. To facilitate direct policy transfer from human interaction data to robotic systems, we employ a segmentation model to filter out embodiment-specific information. Additionally, we leverage task-space tool-action representations to ensure robustness to variations in robot base configurations.
Our contributions are as follows:
  1. 1.
    We propose a novel framework that leverages two-view human interaction data to enable scalable, intuitive, and cost-effective data collection for training robot policies on complex tool-use tasks.
  2. 2.
    We validate the effectiveness of our approach on diverse challenging real-world tool-use tasks (e.g., nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking). Our method achieves a 71% higher success and 77% less data collection time than diffusion policies trained on SpaceMouse [17] or Gello [74] data, with some tasks solvable only by our method. Our method also outperforms handheld grippers like UMI [16], reducing collection time by 41%.
  3. 3.
    We provide an extensive analysis of our method’s robustness under varying conditions, including changes in camera poses, robot base movements, and human-induced perturbations. Additionally, we conduct ablation studies to evaluate the effects of different design choices on policy performance, including embodiment segmentation, random cropping, and novel view augmentation.
II Related Works
Figure 2: Policy Design. Human interaction data was collected using two RGB cameras and processed through the foundation model MASt3R [34] to generate 3D reconstructions. Using 3D Gaussian splatting, we sampled novel views to augument the dataset. The human hand (embodiment) was segmented out from the images to create embodiment-agnostic observations, which serve as inputs to the policy. To label actions for policy training, a pose estimation model (FoundationPose [71] in this work) was used to extract the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$. A coordinate transformation was then applied to compute the tool’s pose in the task space, $T_{\text{tool}}^{\text{task}}$. Finally, a diffusion policy was implemented as the visuomotor policy to enable effective learning and execution.
II-A Data Collection for Robot Learning
High-quality data is essential for training robots to learn and generalize across tasks. Simulation data has been widely used for its cost-effectiveness and scalability [86, 44, 28, 81], but the gap between simulation and real-world performance remains a persistent challenge. To overcome this, many researchers have turned to real-world teleoperated demonstrations, which offer more reliable transferability by minimizing the domain gap between training and testing environments [88, 14, 15, 56, 73]. Advancements in leader-follower devices, like ALOHA [85, 22] and GELLO [74], have simplified robot demonstration data collection but remain tied to specific robot platforms. More recently, portable tools such as hand-held grippers, e.g., UMI [16] and LEGATO [57], have emerged as a promising alternative for in-the-wild data collection. Yet, their high cost and the need for custom robot modifications continue to limit widespread adoption. Tool-based policy representations have emerged as an effective way to collect data for robot learning. MimicTouch uses tactile-based tools for contact-rich manipulation [79], while ScrewMimic models bimanual tasks as constrained screw motions for learning from human videos [3]. However, tactile methods require extra hardware, and the screw motion assumption may not hold. Another line of work by Wen et al. [70] aims to learn category-level representations from a single demo to transfer pose trajectories across similar objects [70], but assumes a static target (e.g., battery slot), limiting real-world applicability in real-world tasks where the spatial configuration of objects may change. Unlike these methods, our method only requires natural human data, without assuming access to tactile sensors or requiring constrained screw motion models, making it significantly more cost-effective, scalable, and accessible.
II-B Cross-Embodiment Policy Learning
Cross-embodiment policy learning enables robots to transfer policies across embodiments, such as those with different kinematic structures [25, 52, 77]. Prior work has explored conditioning policies on embodiment representations using multi-embodiment datasets [18, 69, 80, 13, 24, 52, 2], but these approaches often face challenges in effectively leveraging human interaction data. Recent approaches utilize human data, such as estimating point flow from human video [72] or generating latent plan or high-level plans from human data [37, 68]. However, their dependence on robot data for low-level control limits scalability when robot data is expensive or difficult to collect. Additionally, prior works highlight the importance of masking human and robot embodiments for visual consistency [4, 31]. However, Bahl et al. [4] relies on predefined motion primitives, while Kareer et al. [31] still requires robot data with human data as augmentation. Our approach adopts a similar masking idea but enables robots to learn freely, even agile motions from human videos, without any robot data. Another line of work tokenizes observation inputs and action outputs into a unified transformer network, enabling generalized policy learning across embodiments [59, 76]. However, these methods require large models and extensive datasets, making them resource-intensive and time-consuming to train, and lack the capability for direct policy transfer between embodiments. Overall, reliance on robot-specific hardware and data restricts scalability and accessibility. In contrast, our method leverages natural human data, eliminating the need for robot data as a training source.
II-C Cross-Viewpoint Policy Learning
Viewpoints can change as robots manipulate or interact with their environment. Existing approaches address viewpoint changes by learning view-invariant latent representations [58, 11], leveraging equivariant 3D representations [63, 23, 89], or augmenting input data with diverse viewpoints [51, 87]. An early attempt by Sharma et al. [61] introduced a hierarchical model that generates first-person sub-goals from third-person demonstrations via a high-level module, while a low-level embodiment-specific module predicts actions. Recent work by Chen et al. [12] leveraged the SAM model [32] and ControlNet [83] to transform the viewpoint of one robot into another, performing view synthesis with ZeroNVS [54] to augment datasets and enhance policy learning across robots and viewpoints. Another approach by Yuan et al. [82] combined reinforcement learning with multi-view representation learning and a Spatial Transformer Network (STN) to improve policy robustness in visually complex environments. However, these methods face limitations such as scalability issues in hierarchical models, the complexity of controlled image editing, high computational costs in view synthesis, and the need to build sophisticated simulators for reinforcement learning when adapting to diverse environments. Compared to previous work, our method enables efficient and scalable cross-viewpoint transfer from human data by leveraging two RGB images with the MASt3R [34] foundation model for 3D reconstruction and employing Gaussian splatting for fast and cost-effective novel view synthesis.
III Problem Statement
We formulate the robotic manipulation task as a Markov Decision Process, where the goal is to learn a policy $\pi:\mathcal{O}^{r}\to\mathcal{A}$ that enables a robot to perform a given task. The robot’s observation space, $\mathcal{O}^{r}$, consists of a single-view RGB image $I^{r}\in\mathbb{R}^{128\times 128\times 3}$ and proprioception data $x^{r}\in SE(3)$. To train the policy, we use an imitation dataset composed of $N$ human play, denoted as $D=\{(O_{0}^{h},O_{1}^{h},\dots)\}_{n=1}^{N}$. Each human play observation $\mathcal{O}^{h}$ consists of two RGB images captured from different viewpoints: $\mathcal{O}^{h}=\{I^{v1},I^{v2}\}$, where $I^{vi}\in\mathbb{R}^{480\times 640\times 3}$. We preprocess the dataset to get the action by using a 6D pose estimation and tracking model, resulting in $D=\{(O_{0}^{h},a_{0},O_{1}^{h},a_{1},\dots)\}_{n=1}^{N}$, where each action $a\in SE(3)$.
Our objective is to train a robot to perform the same task as demonstrated in the human dataset while overcoming the embodiment differences. We assume that the tool is rigid, the tool and the end effector are rigidly attached during deployment, and the transformation between the tool and the end effector is estimated once before deployment. Our approach aims to allows the robot generalize task execution across embodiments while retaining the core skills demonstrated by humans.
IV Our Robot Policy Learning Framework
Our framework enables the direct transfer of human play data into deployable robot policies. It is designed to fulfill the following key objectives:
  • Support for Dynamic and High-Precision Tasks: Human play, with its inherent fluidity, enables the execution of highly dynamic tasks. Examples include flipping an egg in a pan or performing other actions that require swift, accurate, and natural motions — challenges that are often difficult to address with traditional teleoperation systems or handheld grippers.
  • Robustness: The framework ensures robust performance under dynamic conditions, enabling reliable task execution even with moving or shaking cameras. While broader deployment on mobile platforms such as quadrupeds or humanoids remains an open challenge, our design and experimental results suggest strong potential for generalization to dynamic environments.
  • Generalization Across Robotic Embodiments and Object Categories: The framework demonstrates broad generalizability, validated on robotic platforms such as the UR5e and Kinova Gen3. It extends its capabilities to manipulate a wide range of object categories, showcasing its adaptability to various tasks, setups, and environments.
  • Affordability and Accessibility: The framework requires only two monocular RGB cameras, such as smartphones, webcams, or RealSense cameras. With approximately 7.14 billion smartphones worldwide — covering around 90% of the global population — this setup is accessible to almost anyone [30]. By relying solely on RGB cameras, the framework eliminates the need for designing, printing, or manufacturing additional hardware during the data collection, ensuring a cost-effective and inclusive solution.
  • Intuitive and Natural Interaction: Users can interact naturally, without the need for specialized equipment or additional tools. Using their bare hands and common tools, participants can intuitively perform a variety of tasks. Our approach removes technical barriers associated with 3D printing and other hardware setups, fostering a seamless, user-friendly experience for data collection.
The following sections outline the framework’s design (Figure 2), its underlying principles, and logistical considerations for development and deployment.
IV-A Tool Usage for Data Collection and Manipulation
Humans naturally and intuitively use tools for everyday tasks such as cooking, eating, cleaning, and interacting with the world. Tools act as extensions of human actions, enabling diverse interaction with objects. The natural relationship between tools and objects provides an ideal interface for training robots to mimic human actions using tools, with minimal gap between human and robot tool usage. While grasping and pick-and-place tasks have been extensively studied in previous works, our work focuses on enabling robots to use the same tools humans commonly employ to interact with their environment effectively with the following benefits:
  • Minimized Embodiment Gap: Abstracting actions to the tool pose reduces morphological dependency, enabling policies to generalize across embodiments.
  • Scalable Data Collection: Simplifies the data collection process by eliminating the need for costly robot-specific demonstrations, making our method more accessible.
During data collection, humans can naturally use tools with their hands without requiring additional devices. For robot deployment, the tool can be attached to the robot in two ways:
  • Rigid Grasping: Grasping or picking, which has been extensively studied in prior works, is demonstrated in our Kinova Gen3 robot experiments and involves the robot securely grasping the tool.
  • Customized Fast Tool Changer: Designed for versatility, the tool changer is compatible with any robot that uses the ISO 9409-1-50-4-M6 flange, as demonstrated in our UR5e experiments.
IV-B Embodiment-Agnostic Perception
To encourage cross-embodiment transfer, we adopt a strategy that reduces the perception gap between the training and deployment phases. During training, human play data is collected, featuring human hands interacting with tools and objects. In deployment, robots execute the learned tasks. As showcased in our experiments, the visual differences between human hands and robotic end-effectors can introduce discrepancies that hinder generalization. To address this, we employ Grounded-SAM [49] to segment and mask out the embodiments in each phase. During training, human hands are masked, while during deployment, the robotic embodiments are masked, which ensures that the remaining parts of the scene in both training and testing phases appear visually similar. By aligning perception across embodiments, the framework mitigates distractions caused by embodiment-specific features, enabling better generalization to human-to-robot policy transfer.
IV-C View Augmentation
We use cameras for data collection due to their availability. With approximately 7.14 billion smartphones equipped with cameras, our approach can scale effectively [30]. However, using data from a single camera introduces challenges such as a lack of 3D perception and sensitivity to camera pose.
3D Reconstruction
To address these issues, we use MASt3R [34], an image-matching model that reconstructs accurate 3D environments from two RGB images. This eliminates the need for additional depth sensors, which are less common and consume more power compared to RGB sensors. We use two cameras, which is the minimum requirement for 3D reconstruction, to capture demonstration data. Using only one camera can lead to scale ambiguity in monocular settings, where poses can be scaled by an arbitrary, scene-dependent factor. By capturing images from two viewpoints, MASt3R reconstructs a 3D point cloud without requiring camera extrinsics or intrinsics, and globally aligns point maps within a multi-view 3D reconstruction framework. The process results in high-quality 3D representations.
Data Augmentation
Using 3D Gaussian splatting, we model the scene and synthesize novel viewpoints from human interaction data, which effectively augments the dataset, generating additional perspectives even if the training data was captured from only two views. These synthetic viewpoints provide the robot with a multi-angle understanding of the scene, allowing the policy to be trained on a more diverse and comprehensive set of visual inputs. Additionally, we apply random cropping to the images for data augmentation before feeding them into the policy network, following the approach from diffusion policy [14, 15]. Random cropping further improves the method’s robustness, enabling the policy to generalize better to variations in visual inputs.
IV-D Action Representation for Tool Manipulation
To support general tool usage, we propose a task-frame, tool-centric action representation denoted as $T_{\text{tool}}^{\text{task}}$. This representation focuses on the tool being manipulated, independent of human or robot morphology or camera pose. A visual depiction of the coordinate systems is provided in Figure 3.
Using a 6D pose estimation model (e.g., FoundationPose [71]), we determine the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$. To make the policy robust to camera and base movement, we transform this into the task frame:
$T_{\text{tool}}^{\text{task}}=T_{\text{camera}}^{\text{task}}T_{\text{tool}}^{\text{camera}},$
where $T_{\text{camera}}^{\text{task}}$ represents the transformation from the camera to the task frame.
TABLE I: Benchmark Attributes of Real-World Tasks. These benchmarks evaluate the precision, adaptability, and capability of our framework to address tasks requiring high precision, handling extreme dynamics, utilizing extrinsic dexterity, performing in contact-rich scenarios, and overcoming gravity.
Benchmark High-Precision Extreme Dynamics Using Extrinsic Dexterity Contact-Rich Overcoming Gravity
Task 1: Nail Hammering
Task 2: Meatball Scooping
Task 3: Pan Flipping (Egg, Bun, Patty)
Task 4: Wine Balancing
Task 5: Soccer Ball Kicking
IV-E Training and Deploying Robot Policies
We use diffusion policy [14] as our policy representation to predict $T_{\text{tool}}^{\text{task}}$, trained using the ACCESS system [7]. During deployment, for stationary robots, the task frame aligns with the base frame. For robots with moving base, base movement is compensated using $T_{\text{task}}^{\text{base}}$. The final end-effector pose in the robot base frame, used as the control command for the robot controller, is computed as:
$T_{\text{eef}}^{\text{base}}=T_{\text{task}}^{\text{base}}T_{\text{tool}}^{\text{task}}T_{\text{eef}}^{\text{tool}},$
where $T_{\text{eef}}^{\text{tool}}$ is the fixed transformation between the tool and the robot end-effector.
Figure 3: Coordinate System Diagram. This diagram illustrates the Camera frame, Tool frame, Task space frame, End-Effector (EEF) frame, and Base frame. We use $T_{\text{tool}}^{\text{task}}$ as the action representation for the visuomotor policy output.
V Policy Evaluations
Our experimental evaluations aim to assess the effectiveness of our framework for deploying robot policy learning across three key dimensions:
  • Capabilities: What skills can our framework enable robots to acquire, and how robust are the policies under challenging conditions like a moving camera or base?
  • Effectiveness: How well does our framework achieve its objectives outlined in Section IV? Can it support reliable and scalable policy learning for complex tasks while streamlining data collection for highly dynamic, contact-rich, or dexterous scenarios?
  • Policy Execution Efficiency: How efficiently do the trained policies execute tasks? Does our framework enable smoother and more natural motion trajectories? Can it achieve faster and more fluid task completion compared to baseline methods?
To evaluate our framework for learning from direct human play, we developed a set of real-world robotic task domains using two embodiments: Kinova Gen3 and UR5e. These tasks are designed to test various aspects of policy capabilities, efficiency and effectiveness. Table I summarizes the key characteristics of these tasks. Visual input for the policies is provided by two RealSense D415 cameras. We describe the tasks in detail, highlighting their challenges and the specific capabilities we aim to test.
Figure 4: Policy Rollouts. We evaluate diverse real-world tasks: nail hammering (precision in locating a nail tip), meatball scooping (slippery object, constrained environments), pan flipping (extremely dynamic, high-speed, contact-rich), wine balancing (precise control of unstable objects), and soccer ball kicking (dynamic object handling, goal-directed actions).
Nail Hammering: The task involves hammering a 3D-printed nail, requiring the robot to locate the nail, draw back the hammer, and strike the nail tip accurately. With a diameter of less than 15.5 mm, the nail tip demands high precision. Challenges include localizing the nail tip precisely and planning effective hammer trajectories. To evaluate generalization, the initial position of the nail is varied across different spatial configurations. We collected 180 seconds of data (40 human play episodes) from a single participant.
Meatball Scooping: In this task, the robot must use a spoon to scoop a meatball from a pan and transfer it to a bowl. This task is challenging due to the complex dynamics of the meatball, which can roll unpredictably within the pan. Additionally, the interaction between the spoon and the meatball requires careful control, as improper contact can cause the meatball to slip or escape the spoon. We randomize the initial position of the meatball within the pan to test its generalization capability. We collected 340 seconds of data (50 human play episodes) from a single participant.
Pan Flipping (Egg, Burger Bun, Meat Patty): The objective of this task is to use a pan to flip various objects, such as an egg, a burger bun, and a meat patty. The task is challenging due to its high-speed dynamics, requiring the robot to overcome gravity and accurately manage the interaction between the pan and the objects. Each object differs in weight, shape, and texture, adding further complexity. This task evaluates the policy’s ability to handle fast, contact-rich interactions and adapt to diverse object types. To increase variability, the initial positions of the objects within the pan are randomized. Furthermore, the rapid and dynamic nature of the task makes it unsuitable for classical demonstration collection methods, highlighting the advantages of using bare-handed human play for data collection. We collected 50 seconds of data (38 human play episodes) from a single participant using three different pans and two object types.
Wine Balancing: In this task, the robot needs to use a hook to lift a wine bottle and carefully insert it into an unstable, zero-gravity wine rack. The task is challenging due to the precise control required to suspend the bottle in mid-air and counteract gravitational forces effectively. Any over-insertion or under-insertion will cause the bottle to lose balance. To constrain the horizontal movement of the rack, screws were added as obstacles to limit lateral motion. No additional variability was introduced. We collected 223 seconds of data (15 human play episodes) from a single participant.
Soccer Ball Kicking: In this task, the robot must use a golf club to kick a ball that slides into a field and direct it into the goal. To increase the challenge, a 3D-printed row of players serves as obstacles between the robot and the goal. The task is difficult because the robot must accurately intercept the moving ball, strike it with the correct force and direction, and ensure it avoids obstacles before reaching the goal. The position of the player obstacle varies. We collected 78 seconds of data (20 human play episodes) from a single participant.
Baselines: The primary focus of this work is to demonstrate the effectiveness and efficiency of learning directly from human play without relying on robot-generated data. We compare our approach against two baselines: a diffusion policy trained on robot demonstrations, and UMI [16], a hand-held gripper-based data collection method. The robot demonstration dataset is collected using either a SpaceMouse or Gello interface under the same data collection time. Additionally, we perform ablation studies to analyze the impact of key components, such as random cropping of images before policy training, novel view synthesis-based data augmentation, and the effects of embodiment segmentation. To further illustrate the advantages of our approach, we compare trajectory rollouts for a meatball-scooping episode, highlighting how our method is more sample-efficient and less prone to distribution shifts by eliminating excessive waypoints.
Evaluation Metrics: During testing, we introduce two types of variations: (1) randomizing the initial spatial configurations of objects in each task to assess policy generalization, and (2) varying camera positions to evaluate the robustness of policies to different viewpoints. All methods, including the baseline and ablation variants, are tested under the same conditions. Performance is evaluated using two metrics: success rate, which measures the proportion of successfully completed task trials and reflects policy effectiveness, and task completion time, which captures the average duration to complete tasks and reflects policy efficiency.
TABLE II: Task Success Rates and Completion Times. Success rates are the number of successful trials out of total episodes, and average completion times are based on successful trials. “DP” refers to the diffusion policy trained on teleoperation data. “Not Feasible” tasks denote cases where teleoperation failed due to extreme dynamics, precision, or reactivity demands. Our method consistently achieves higher success rates and shorter completion times.
Task Method Success Rate Time (s)
Hammer Nailing DP 0/13 -
Ours 13/13 11.0
Meatball Scooping DP 5/12 42.0
Ours 10/12 12.4
Pan Flipping - Egg DP Not Feasible -
Ours 12/12 1.5
Pan Flipping - Burger Bun DP Not Feasible -
Ours 9/12 1.9
Pan Flipping - Meat Patty DP Not Feasible -
Ours 10/12 2.3
Wine Balancing DP Not Feasible -
Ours 8/10 30.9
Soccer Ball Kicking DP Not Feasible -
Ours 6/10 2.0
TABLE III: Task success rates comparing our method with the hand-held gripper-based method on Nail Hammering.
Method Demo Duration & Count Success Rate
UMI [16] $\sim$180 seconds (25 demos) 0/13
UMI $\sim$720 seconds (100 demos) 13/13
Ours $\sim$180 seconds (40 demos) 13/13
VI Experiment Results
In our experiments, we demonstrate that our framework is both effective and efficient for training robots with advanced capabilities. Furthermore, leveraging human play data enables robots to perform smoother movements and acquire skills that are challenging or even impossible to achieve with robot-generated data.
Figure 5: Policy Testing Across Camera Poses in Nail Hammering. (a) Camera poses for data collection and evaluation. (b-d) Performance ranges for methods trained with/without random cropping (RC) and view augmentation (VA).
Figure 6: Policy Execution Trajectory Comparison. (a) Initial setup for meatball scooping. (b) Comparison of end-effector XY trajectories from our framework and a policy trained on robot-collected data. Our method generates smoother, more natural, and efficient trajectories with fewer waypoints, enabling fluid and precise motion. In contrast, the robot-collected policy produces jerky movements with excessive waypoints, which hinder task performance.
VI-A Capabilities and Effectiveness
Table II presents the results of our real-world robot tasks, showing that our framework consistently outperforms baseline methods by achieving significantly higher success rates across all evaluated scenarios. We further compare our method with a stronger hand-held gripper baseline, UMI [16], as shown in Table III. In our default setup, SLAM-based mapping failed due to low environmental texture. To address this limitation, we added a textured background to support reliable mapping for UMI. For the nail hammering task, we evaluated UMI using 25 demonstrations (matching our collection time) and 100 demonstrations (to assess its ideal performance). UMI failed all 13 trials with 25 demonstrations but succeeded in all 13 trials with 100. UMI could not be applied to the wine balancing task due to contact-induced tool displacement, nor to the pan flipping task due to tool inertial slippage. In the soccer kicking task, large and fast motions made it nearly impossible to localize the demonstration trajectory within the initial map.
In contrast to the baselines, our method demonstrates reliable performance across all tasks. As illustrated in Figure 4, our method excels in real-world policy rollouts by accurately detecting spatial locations in tasks such as nail hammering and meatball scooping. For pan flipping tasks, it performs the high-speed motions required to effectively flip eggs, burger buns, and meat patties. Additionally, it demonstrates precise control in lifting and inserting a wine bottle into its stand and reacts swiftly and accurately to kick a soccer ball in the soccer ball kicking task. The superior performance of our framework is primarily due to its ability to collect a significantly larger volume of episodes within the same data collection timeframe, as detailed later in Section VII. This broader dataset covers a wider range of task variations, enabling more robust and adaptable policy training. Furthermore, our approach overcomes the inherent limitations of Gello and SpaceMouse, enabling the collection of demonstrations for scenarios that these tools cannot adequately handle due to their constraints.
VI-B Policy Execution Efficiency
Our framework demonstrates exceptional efficiency in task execution, achieving faster task completion times and producing smoother action motions compared to baseline methods, as shown in Table II. The efficiency is largely attributed to the nature of human play data, which captures the fluidity and speed of real-world human activities, resulting in smoother and more natural trajectories in the training dataset. In contrast, previous approaches relying on teleoperated data often suffer from significantly slower speeds and less natural motions, limiting their effectiveness in dynamic scenarios. By leveraging the realistic dynamics of human play, our framework not only accelerates task execution but also enhances motion quality, making it better suited for real-world applications.
Figure 7: Initial States for All Evaluation Episodes. All methods are evaluated using the same set of manually defined initial states, overlayed in the image. These states ensure diverse variations to test the policy’s spatial generalization capabilities.
Figure 8: Robustness to Camera and Base Movement. (a) Camera Pose Robustness: The policy demonstrated the ability to handle camera shaking across three tasks—meatball scooping, nail hammering, and pan flipping. The first row shows the camera view, while the second row provides a scene overview with the shaking motion. (b) Robot Base Robustness: The policy successfully compensated for base shaking, even when the shaking frequency exceeded the robot’s control frequency. (c) Chicken Head Stabilization: At lower base movement frequencies, the end effector displayed a stabilization effect similar to a chicken’s steady head. (d) Combined Robustness: The policy maintained task performance under simultaneous camera and base shaking.
Figure 9: Human Perturbation Robustness. This figure showcases the robot’s ability to handle human-induced perturbations across three tasks: (1) In nail hammering, the robot successfully followed a manually moved nail; (2) In meatball scooping, it located and scooped meatballs even when additional ones were thrown into the pan mid-task; and (3) In egg flipping, the robot consistently flipped the egg back after human intervention repositioned it. These results highlight the policy’s robustness to unpredictable human perturbations.
VI-C Benefits of Tool-Based Action Representation in Task Space
We observed that using the tool in the camera frame as the action representation allows the policy to perform comparably to prior works when the camera is stationary. However, the success rate drops to zero when the camera is moving. Tracking the camera pose in real time during motion is non-trivial, leading to incorrect calculations of the end-effector position in the base frame. When the base moves, policies relying on the tool in the base frame fail entirely, achieving zero success. These failures occur because such policies assume a fixed relative transformation between the base and the workspace, an assumption invalidated by base motion. In contrast, representing actions using the tool in task space enables the policy to remain effective under base perturbations. Our approach demonstrates robust performance when camera poses vary, when the base moves, and even during simultaneous shaking of the camera and base.
VI-D Effects of Random Cropping and View Augmentation
Our experiments show that random cropping (RC) and view augmentation (VA) enhance policy robustness to diverse camera poses. Training with RC improves resilience to minor camera perturbations, such as small movements or shaking, while VA extends robustness by exposing the model to a broader range of camera configurations during training. We evaluated these techniques on the nail hammering task and present the results in Figure 5, comparing three models: one trained with both RC and VA, one with RC only, and one without either technique. The combined use of RC and VA significantly expands the working range of camera configurations, including those with substantial deviations from the original viewpoint. The robustness arises from the model’s exposure to a diverse set of camera poses during training.
VI-E Policy Execution Trajectory Comparison
Our framework generates faster, smoother, and more natural trajectories compared to traditional approaches, as demonstrated by the end effector (EEF) XY trajectory for the meatball scooping task in Figure 6. Figure 6(a) depicts the initial setup of one trial, while Figure 6(b) contrasts the XY trajectory of the policy rollout using our framework with that of a policy trained on robot-collected data. Our framework surpasses the baseline approach in two critical aspects:
  • Improved Motion Fluidity: Trajectories generated by our framework are significantly smoother and less jerky, enabling more fluid and precise task execution. In contrast, the robot-collected data results in jerky motions caused by excessive waypoints, which disrupt movement precision and degrade performance.
  • Efficient and Effective Learning: The trajectories produced by our framework contain 10 times fewer waypoints than those from robot-collected data. The reduction simplifies the learning space, mitigates cumulative error growth, and minimizes the risk of distribution shifts commonly observed in behavior cloning. As a result, our policies exhibit significantly improved sample efficiency.
VI-F Generalization
Spatial Generalization: Our policy demonstrates strong generalization across various spatial configurations. Figure 7 illustrates the range of initial states tested for multiple tasks. In the hammer nailing task, nail positions vary. For meatball scooping, meatball positions are tested. In soccer ball kicking, we evaluate two goalkeeper configurations. For pan flipping, object poses vary to test all areas of the pan. As shown in Table II, our method achieves high success rates across these ranges.
Object Generalization: Our method generalizes effectively to different objects in pan flipping tasks, including a toy egg, a 3D-printed meat patty, and a real burger bun (Figure 7, second column). With only 13 demonstrations, the policy succeeds by leveraging a simple but effective strategy: tilting the pan to slide the object into a corner, then flicking the pan to propel and flip the object. Our manipulation approach for pan flipping enables robust generalization across diverse object types.
Tool Generalization: To assess the generalization ability of our policy across different tools, we conducted a pan-flipping experiment using a burger bun and five pans: large, medium, small, tiny, and square. For each pan, we collected 12 trials with varying initial configurations and reported the success rates (Fig. 10). The policy was trained on demonstrations using the large, medium, and square pans, and evaluated on all five. Results indicate that our method exhibits some generalization across both pan sizes and shapes (circular vs. square). High success rates were observed with the large and medium pans. However, performance declined on smaller pans, likely due to their limited surface area. The square pan also showed lower success rates, as its shallow edges allowed the bun to slide out during flipping.
Figure 10: Tool Generalization (a) The tested pans. (b) Success rate across 12 testing trials.
TABLE IV: Comparison of Data Collection Methods. This table compares various data collection methods for robotics. For cost, we calculate only the additional expenses required for data collection, excluding cameras, as they are considered a basic and commonly used sensor for robots rather than an additional purchase. Each method is assessed based on cost, ease of use, required expertise, precision, and maintenance effort. Our method stands out as cost-free, easy to use, highly precise, and requiring minimal maintenance.
Method Cost Ready-to-Use Pre-Knowledge Required Precise Maintenance Expense
Visual Imitation Made Easy [78] \$340 No Yes No Moderate
UMI [16] \$371 No Yes Yes Moderate
LEGATO [57] \$1060 No Yes Yes Moderate
Spacemouse [17] \$169 Yes Yes Yes Low
VR (Meta Quest 2 [40]) \$300 Yes Yes No Moderate
AR (Apple Vision Pro [1]) \$3499 Yes Yes Yes High
Gello [74] \$272 No Yes No Moderate
Ours \$0 Yes No Yes Minimal
Figure 11: Data Collection Efficiency and Reliability. (a) Human hands excel in manipulation tasks, leveraging natural and intuitive efficiency. (b) Failure cases for Gello and Spacemouse include insufficient speed, lack of tactile feedback during data collection, safety stops, collisions, teleoperation delays, and difficulty handling high-speed or complex tasks. (c) Failure cases for handheld grippers such as UMI [16], where issues arise from tool slippage due to inertia or displacement caused by contact forces.
Figure 12: Quantitative Comparison of Data Collection Methods. Human hands significantly reduce data collection time—by 73% for nail hammering and 81% for meatball scooping—while maintaining consistently low variation across trials. Teleoperation fails entirely in dynamic tasks and tasks requiring high precision. In nail hammering, human hands are 41% faster than UMI [16], which also struggles in dynamic scenarios and low-texture environments.
VI-G Robustness
Camera Pose Robustness: We evaluated the policy’s ability to handle camera pose variations by introducing camera shaking in three tasks: meatball scooping, nail hammering, and pan flipping (egg), as shown in Figure 8(a). The first row shows the camera view, while the second row shows the scene overview and the shaking motion. Despite the disturbances, the policy successfully completed all tasks. This robustness stems from random cropping during training, which allows the policy to adapt to partially cropped inputs and minor visual changes.
Robot Base Robustness: To evaluate robustness to base movement, we manually shook the robot base during task execution (Figure 8(b)). When the shaking frequency exceeded the control frequency, the end effector oscillated with the base. However, the task-space action design enabled the robot to compensate for these movements and maintain task success. Additionally, as shown in Figure 8(d), the policy remained effective even when both camera and base shaking were applied simultaneously.
Chicken Head Stabilization: At lower frequencies of base movement, where the shaking was slower than the robot’s control frequency, the end effector exhibited a stabilization effect similar to a chicken’s head remaining steady [75], as shown in Figure 8(c). The behavior highlights the robot’s ability to maintain precise control during mild perturbations.
Human Perturbation Robustness: We further tested the policy’s resilience to human-induced perturbations across three tasks, as shown in Figure 9. In the nail hammering task, the robot successfully tracked and followed a manually moved nail by a human. During meatball scooping, the robot located and scooped meatballs, even when additional meatballs were thrown into the pan mid-task. For egg flipping, the robot consistently flipped the egg back each time a human intervened and returned the egg to different initial positions. These results underscore the policy’s robustness in maintaining reliable performance under real-time human interactions.
VII Data Collection Efficiency and Affordability
We compare various data collection methods for robot imitation learning, focusing on throughput, reliability, cost, usability, and precision. Our evaluation includes teleoperation tools like Gello and Spacemouse for 6DOF (UR5e) and 7DOF (Kinova Gen3) robots, alongside methods such as Visual Imitation Made Easy, handheld grippers (e.g., UMI and LEGATO), and devices like VR (Meta Quest 2), AR (Apple Vision Pro), and Kinematic replicate (Gello).
VII-A Data Collection Efficiency
Our framework achieves significantly higher data collection throughput than traditional methods, enabling more demonstrations within the same timeframe. The improvement is driven by the natural and intuitive efficiency of human play, which ensures faster and more reliable task execution. Figure 11(a) highlights the superior manipulation capabilities of human hands, while Figure 12 quantifies the substantial time savings per episode. For nail hammering and meatball scooping, Gello and Spacemouse were used as teleoperation methods, respectively. Human hands reduced data collection time by 73% and 81% for nail hammering and meatball scooping, with consistently low variation in performance. In more complex tasks like pan flipping, wine balancing, and soccer ball kicking, teleoperation methods failed entirely due to limitations such as lack of tactile feedback, delays, and difficulty handling dynamic or precise actions. Our method further reduces data collection time by 41% compared to handheld grippers such as UMI [16] in nail hammering. UMI proved ineffective in wine balancing and pan flipping due to tool inertial slippage or contact-induced displacement, and failed in soccer kicking because of difficulty localizing large, fast motions. Moreover, it requires rich textures to build a pre-collection map, which our method does not. These results underscore the superior efficiency, robustness, and versatility of human play as a scalable solution for high-quality robot learning datasets.
VII-B Reliability
Figure 11(b) and Figure 11(c) illustrates typical failure cases with Gello, Spacemouse, and UMI [16], which frequently encounter issues such as safety stops or collisions during data collection. In contrast, our method ensures smooth, uninterrupted operation, avoiding these limitations. Traditional methods face significant challenges in high-speed or complex tasks. For example, Gello and Spacemouse struggle with replicating the extreme dynamics and precise motions required for flipping objects like eggs during pan flipping, often resulting in unsuccessful attempts. Similarly, teleoperation delays prevent timely strikes during soccer ball kicking, consistently leading to missed kicks and repeated failures. In tasks like wine balancing, the absence of tactile feedback impairs precision during the data collection, causing the wine bottle to tip over during data collection. Furthermore, in meatball scooping, the velocity vectors generated by Spacemouse input lead to jerky trajectories with redundant waypoints, significantly reducing efficiency. These challenges make effective training impractical with traditional methods. By leveraging human play, our framework not only addresses these limitations but also provides a reliable and scalable solution for dynamic and precision-demanding tasks.
VII-C Discussion of Data Collection Methods
Table IV compares various data collection methods based on cost, usability, expertise requirements, intuitiveness, and precision. Our method incurs no additional cost (\$0), unlike hardware-dependent solutions like UMI and LEGATO, which demand significant investment. This affordability makes our approach accessible to users from diverse backgrounds without financial constraints. Unlike hardware-based systems such as UMI, LEGATO, Gello, and Spacemouse, which are prone to malfunctions and maintenance issues, our hardware-free framework ensures reliability and eliminates repair delays or expenses. Additionally, it requires no supplementary 3D printing, in contrast to approaches like Visual Imitation Made Easy, UMI, and LEGATO. The simplicity of our design promotes inclusivity in collecting large-scale dataset for robot learning research. Our method also offers a more natural experience compared to tools like Spacemouse, while being far more cost-effective than VR and AR devices. Moreover, systems like Gello and Spacemouse lack the precision necessary for dynamic tasks, a limitation addressed by our approach. Overall, our method is a cost-effective, and accessible solution for data collection, overcoming key drawbacks of existing approaches while reducing complexity and maintenance needs.
VIII Limitations and Future Work
Our framework has certain limitations. First, the perception pipeline relies on FoundationPose for extracting the tool’s pose during manipulation. Errors in pose estimation may occasionally require data recollection, adding time and effort. Improving the reliability of the perception pipeline through more robust pose estimation algorithms or self-correction mechanisms is a promising direction for future work. Second, for novel view augmentation, significant noise and reduced realism are observed when augmented views deviate too far from the collected camera views, which can hinder policy performance. Future efforts could focus on leveraging advanced rendering techniques to enhance the realism of augmented views and improve policy generalization. Third, we assume the tool is rigidly attached to the robot’s end effector; however, in real-world, contact-rich manipulation, minor shifts may occur, potentially affecting performance. Addressing this issue by incorporating tactile sensing could improve performance in contact-intensive tasks. Additionally, our method assumes a rigid tool and does not account for flexible or soft tools. Future work could explore using flexible representations for tool state estimation to better handle deformable tools in real-world manipulation scenarios.
IX Conclusion
In this work, we presented a novel framework for human-to-robot imitation learning that leverages human play data to bridge the embodiment gap and enables robust policy training for diverse tool-use tasks. Unlike traditional data collection methods, which are often costly, hardware-dependent, and require technical expertise, our framework democratizes data collection by removing the need for specialized equipment or prior knowledge. Our approach makes data collection more accessible and scalable, empowering broader adoption in robotic learning. We validated our framework across a range of challenging tasks, including nail hammering, meatball scooping, pan flipping with various objects, wine bottle balancing, and soccer ball kicking. The results demonstrate the framework’s superior performance, robustness to variations in camera poses and base movements, and adaptability to different embodiments, such as 6-DOF and 7-DOF robots. By enhancing accessibility, scalability, and reliability, our work lays a strong foundation for advancing robotic manipulation in complex, real-world scenarios.
Acknowledgments
We thank Professor Wenzhen Yuan, Xiaoyu Zhang, Yucheng Mo, Yilong Niu, Hyoungju Lim, and Amin Mirzaee for their support and assistance in reproducing the UMI. We thank Professor Saurabh Gupta, Professor Junyi Geng, Shuijing Liu, Lujie Yang, Neeloy Chakraborty, Hongkai Dai, Jacob Wagner, Shaoxiong Yao, Wei-Cheng Huang for their insightful feedback and suggestions. This work was supported by ZJU-UIUC Joint Research Center Project No. DREMES 202003, funded by Zhejiang University. This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. Additionally, this work used NCSA Delta GPU at NCSA through allocation CIS240753 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Appendix
A-A Implementation Details
A-A1 Hardware Design
We designed two fast tool changers compatible with robots using the ISO 9409-1-50-4-M6 flange, as shown in Figure 13. The left design utilizes a screw mechanism to accommodate general tools, while the right design employs clips for tools with specific mounting shapes.
Figure 13: Fast Tool Changer. Two designs are shown: the left accommodates general tools with a screw mechanism, and the right clips onto tools with specific mounting shapes.
A-A2 Tool Pose Estimation
We use Polycam to scan the tool and obtain its mesh. The mesh is later feed into FoundationPose [71] for 6D pose estimation.
A-B Additional Results
A-B1 Effects of Embodiment Segmentation
Embodiment Segmentation masks the agent’s embodiments during data collection and policy deployment, ensuring visually consistent scenes and reducing the training-deployment visual gap. Embodiment Segmentation significantly improves policy performance, as shown in Figure 14. Figure 14(a) highlights failure cases without segmentation. In the wine balancing task, the robot strikes the table, triggering safety stops due to improper bottle handling. In the soccer ball kicking task, the robot’s actions are inconsistent, shorter, and less precise than during training. Quantitative results in Figure 14(b) further underscore segmentation’s impact. Across 10 trials, segmentation enabled 8 successes in the wine balancing task, while the model without it achieved none. Similarly, in the soccer ball kicking task, segmentation resulted in 6 successes, compared to 2 without it. By aligning training and testing visual distributions, Embodiment Segmentation ensures consistent and reliable robot performance during the training and deployment.
Figure 14: Effects of Embodiment Segmentation. (a) Failure cases without segmentation: In the wine balancing task, the robot strikes the table, triggering safety stops. In the soccer ball kicking task, it performs shorter, less precise actions. (b) Quantitative results: Segmentation improved success rates in wine balancing (8 vs. 0) and soccer ball kicking (6 vs. 2) by reducing the visual gap between training and deployment.

### v2

@@
-Tool-as-Interface: Learning Robot Policies from Human Tool Usage through Imitation Learning
+Tool-as-Interface: Learning Robot Policies from Observing Human Tool Use
 Abstract
-Tool use is critical for enabling robots to perform complex real-world tasks, and leveraging human tool-use data can be instrumental for teaching robots.
-However, existing data collection methods like teleoperation are slow, prone to control delays, and unsuitable for dynamic tasks.
-In contrast, human natural data—where humans directly perform tasks with tools—offers natural, unstructured interactions that are both efficient and easy to collect.
-Building on the insight that humans and robots can share the same tools, we propose a framework to transfer tool-use knowledge from human data to robots.
-Using two RGB cameras, our method generates 3D reconstruction, applies Gaussian splatting for novel view augmentation, employs segmentation models to extract embodiment-agnostic observations, and leverages task-space tool-action representations to train visuomotor policies.
-We validate our approach on diverse real-world tasks, including meatball scooping, pan flipping, wine bottle balancing, and other complex tasks.
-Our method achieves a 71% higher average success rate compared to diffusion policies trained with teleoperation data and reduces data collection time by 77%, with some tasks solvable only by our framework.
-Compared to hand-held gripper, UMI [16], our method cuts data collection time by 41%.
-Additionally, our method bridges the embodiment gap, improves robustness to variations in camera viewpoints and robot configurations, and generalizes effectively across objects and spatial setups.
+Tool use is essential for enabling robots to perform complex real-world tasks, but learning such skills requires extensive datasets.
+While teleoperation is widely used, it is slow, delay-sensitive, and poorly suited for dynamic tasks.
+In contrast, human videos provide a natural way for data collection without specialized hardware, though they pose challenges on robot learning due to viewpoint variations and embodiment gaps.
+To address these challenges, we propose a framework that transfers tool-use knowledge from humans to robots.
+To improve the policy’s robustness to viewpoint variations, we use two RGB cameras to reconstruct 3D scenes and apply Gaussian splatting for novel view synthesis.
+We reduce the embodiment gap using segmented observations and tool-centric, task-space actions to achieve embodiment-invariant visuomotor policy learning.
+We demonstrate our framework’s effectiveness across a diverse suite of tool-use tasks, where our learned policy shows strong generalization and robustness to human perturbations, camera motion, and robot base movement.
+Our method achieves a 71% improvement in task success over teleoperation-based diffusion policies and dramatically reduces data collection time by 77% and 41% compared to teleoperation and the state-of-the-art interface, respectively.
 Figure 1: Tool-as-Interface.
@@
 The framework enables robots to learn robust policies that can operate effectively under challenging conditions, such as base and camera movement, and achieve high performance on a variety of complex manipulation tasks.
-I Introduction
-Tool use is essential to how humans interact with and transform their environment [67, 41].
-For instance, humans use a pan to fry food and flip it, ensuring even cooking on both sides.
-Despite its significance, tool use beyond parallel jaw grippers remains underexplored in robotics, with research primarily focused on simpler tasks like grasping and pick-and-place operations [6, 46, 10, 36, 35].
-In this paper, we focus on cost-effective data collection and efficient training of robot policies to rapidly acquire tool-use skills.
-Imitation learning provides a promising pathway for robots to acquire tool-use skills by directly learning from human demonstrations [21, 26, 27, 29].
-The paradigm excels in handling diverse tool-use tasks, as it bypasses the need for task-specific programming by relying solely on human demonstrations.
-However, its full potential hinges on addressing key challenges in collecting high-quality training data.
-Various teleoperation systems [33, 56, 8, 20, 29, 50, 84, 68, 39, 66, 45, 9] and hand-held grippers [65, 19, 53, 47, 43] have been developed to facilitate the collection of high-quality data.
-Teleoperation methods, such as kinematic replication and hand or body retargeting, show great potential [85, 22, 74, 62, 48].
-However, their reliance on direct access to robot hardware limits both practicality and scalability.
-Hand-held grippers [57, 16] offer an alternative by enabling demonstrations in diverse environments.
-While they reduce dependency on robotic systems, their high costs and the technical expertise required for tasks like 3D printing and assembly restrict their accessibility to a specialized group of users.
-To address these limitations, we turn to human interaction data — a natural, intuitive method through which humans interact with their environment during everyday activities without relying on external devices or specialized setups.
-Human interaction data refers to the natural process in which humans use their hands to operate tools and interact with and manipulate the environment freely.
-Unlike controlled demonstrations that require expensive hardware or meticulous preparation, human interaction data involves the spontaneous use of tools to interact with the environment.
-Human interaction data is an accessible, scalable, and cost-effective approach to data collection, requiring no prior knowledge or technical expertise, such as 3D printing or assembly, from participants.
-However, existing methods struggle to fully harness the potential of human interaction data.
-Key challenges include the embodiment gap and the reliance on single-view data, which limits the insights that can be drawn from interaction data [60, 4, 38, 5, 64, 42, 55].
-Our framework addresses these challenges by leveraging the observation that humans and robots can share the same tools.
-We propose a novel approach that utilizes human interaction data to train robust and adaptable robot policies for diverse tool-use tasks (Figure 1).
-Our method minimizes reliance on expensive hardware, making data collection more scalable and accessible to non-experts.
-By capturing 3D information using two RGB cameras and generating 3D reconstructions, our method enables view-invariant policy learning through novel view augmentation.
-To facilitate direct policy transfer from human interaction data to robotic systems, we employ a segmentation model to filter out embodiment-specific information.
-Additionally, we leverage task-space tool-action representations to ensure robustness to variations in robot base configurations.
-Our contributions are as follows:
-  1. 1.
-    We propose a novel framework that leverages two-view human interaction data to enable scalable, intuitive, and cost-effective data collection for training robot policies on complex tool-use tasks.
-  2. 2.
-    We validate the effectiveness of our approach on diverse challenging real-world tool-use tasks (e.g., nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking).
-Our method achieves a 71% higher success and 77% less data collection time than diffusion policies trained on SpaceMouse [17] or Gello [74] data, with some tasks solvable only by our method.
-Our method also outperforms handheld grippers like UMI [16], reducing collection time by 41%.
-  3. 3.
-    We provide an extensive analysis of our method’s robustness under varying conditions, including changes in camera poses, robot base movements, and human-induced perturbations.
-Additionally, we conduct ablation studies to evaluate the effects of different design choices on policy performance, including embodiment segmentation, random cropping, and novel view augmentation.
-II Related Works
+Keywords: Tool Use, Data Collection, Learning from Video
+1 Introduction
+Tool use enables humans to perform complex tasks by extending their physical capabilities.
+In contrast, robotic systems remain largely limited to grasping and pick-and-place operations [1, 2, 3, 4, 5, 6, 7].
+To enable richer manipulation skills, robots must learn to use diverse tools in dynamic environments.
+This work focuses on the efficient training of robot policies for tool use, with an emphasis on scalable and low-cost data collection.
+Imitation learning (IL) is a promising way to acquire tool-use skills from human demonstrations [8, 9, 10].
+Prior work has leveraged teleoperation platforms [11, 12, 10] and hand-held grippers [13, 14] to provide precise supervision.
+However, these systems often require expensive hardware, 3D-printed tools, or expert calibration, hindering scalability across diverse users and environments.
+As a scalable alternative, we learn from natural human manipulation videos.
+While abundant, this data introduces challenges from viewpoint differences and human-robot embodiment gaps [15, 16, 17].
+We introduce a new framework that leverages two-view human manipulation videos and appropriate representations of states and actions to train robot policies for tool use.
+To close the viewpoint gap, our system uses 3D scene reconstruction and novel view synthesis to train a viewpoint-invariant policy.
+To bridge the embodiment mismatch, we filter out embodiment-specific features like hands using segmentation, and employ a task-space, tool-centric action representation supports robustness to robot base variation (Figure 1).
+Our contributions are as follows: (1) We introduce a framework for scalable, intuitive, and cost-effective data collection for robot tool-use learning, using two-view human manipulation videos without requiring teleoperation or specialized hardware; (2) We demonstrate strong generalization across diverse real-world tool-use tasks (e.g., nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking) achieving a 71% higher success rate and 77% reduction in data collection time compared to diffusion policies trained on SpaceMouse [18] or Gello [19], and a 41% improvement over handheld grippers like UMI [14]; and (3) We provide a detailed robustness analysis, evaluating performance under changes in viewpoint, robot base configuration, and human motion, along with ablations on segmentation, novel view synthesis, and random cropping.
+2 Related Works
+Data Collection for Robot Learning: High-quality data is essential for teaching robots new skills.
+Simulation offers scalability and low cost [20, 21, 22, 23], but the sim-to-real gap remains challenging.
+Learning from real-world teleoperated demonstrations helps train policies with minimal distribution shift between training and testing [24, 25, 26, 27, 28].
+Leader-follower systems (e.g., ALOHA [29, 30], GELLO [19]) simplify teleoperation by providing intuitive kinematic replication but require real robots and are costly.
+Portable hand-held grippers (e.g., UMI [14], LEGATO [13]) enable flexible data collection but still require specialized hardware.
+Tool-based policy representations offer another direction.
+MimicTouch leverages tactile feedback for contact-rich skills [31], while ScrewMimic models bimanual tasks as constrained screw motions for learning from human videos [32].
+However, tactile methods require extra hardware, and the screw motion assumption may not hold.
+Another approach transfers single-demonstration trajectories across objects but assumes static object configurations [33], which will not work with changing spatial configurations.
+In contrast, our method uses natural human demonstrations without tactile sensors, special tools, or motion constraints—enabling scalable, low-cost data collection in unconstrained environments.
+Cross-Embodiment Policy Learning: Cross-embodiment learning enables policy transfer across robots with diverse kinematic structures [34, 35, 36].
+Prior approaches relying on multi-embodiment datasets [37, 38, 39, 40, 41, 35, 42] focus solely on data from different robot embodiments and struggle to leverage human demonstrations effectively.
+Recent methods tokenize observations and actions into unified transformer networks [43, 44], but require large models, extensive datasets, and cannot directly transfer policies across embodiments.
+Other efforts use human data to estimate point flow [45] or generate high-level plans [46, 11], but still depend on robot-generated data for low-level control.
+Additionally, prior works emphasize visual consistency through embodiment masking [16, 47].
+However, Bahl et al.
+[16] relies on predefined motion primitives and Kareer et al.
+[47] requires robot data for augmentation.
+As a step further, our approach adopts a similar masking idea but enables robots to learn freely, even agile motions from human videos without any robot data.
+Cross-Viewpoint Policy Learning: Robots can encounter viewpoint variations when interacting with their environments.
+Prior work addresses this by learning view-invariant representations [48, 49], equivariant 3D features [50, 51, 52], or augmenting datasets with varied viewpoints [53, 54].
+An early attempt by Sharma et al.
+[55] introduced a hierarchical model mapping third-person demonstrations to first-person sub-goals, with a low-level module predicting actions.
+Recent work by Chen et al.
+[56] leverages SAM [57] and ControlNet [58] to transform robot images across different robots, employing ZeroNVS [59] for novel-view synthesis to augment datasets and improve policy generalization across viewpoints.
+Another approach by Yuan et al.
+[60] integrates reinforcement learning, multi-view representation learning, and a Spatial Transformer Network to enhance policy robustness in visually complex environments.
+Despite promising results, these methods face scalability issues, reliance on computationally intensive view synthesis, or the need for tailored simulators.
+In contrast, our approach achieves efficient and scalable cross-viewpoint transfer directly from human data by leveraging the MASt3R [61] foundation model for 3D reconstruction from two RGB images and employing Gaussian splatting for fast, cost-effective novel-view synthesis.
+3 Tool-as-Interface Framework
 Figure 2: Policy Design.
-Human interaction data was collected using two RGB cameras and processed through the foundation model MASt3R [34] to generate 3D reconstructions.
-Using 3D Gaussian splatting, we sampled novel views to augument the dataset.
-The human hand (embodiment) was segmented out from the images to create embodiment-agnostic observations, which serve as inputs to the policy.
-To label actions for policy training, a pose estimation model (FoundationPose [71] in this work) was used to extract the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$.
-A coordinate transformation was then applied to compute the tool’s pose in the task space, $T_{\text{tool}}^{\text{task}}$.
-Finally, a diffusion policy was implemented as the visuomotor policy to enable effective learning and execution.
-II-A Data Collection for Robot Learning
-High-quality data is essential for training robots to learn and generalize across tasks.
-Simulation data has been widely used for its cost-effectiveness and scalability [86, 44, 28, 81], but the gap between simulation and real-world performance remains a persistent challenge.
-To overcome this, many researchers have turned to real-world teleoperated demonstrations, which offer more reliable transferability by minimizing the domain gap between training and testing environments [88, 14, 15, 56, 73].
-Advancements in leader-follower devices, like ALOHA [85, 22] and GELLO [74], have simplified robot demonstration data collection but remain tied to specific robot platforms.
-More recently, portable tools such as hand-held grippers, e.g., UMI [16] and LEGATO [57], have emerged as a promising alternative for in-the-wild data collection.
-Yet, their high cost and the need for custom robot modifications continue to limit widespread adoption.
-Tool-based policy representations have emerged as an effective way to collect data for robot learning.
-MimicTouch uses tactile-based tools for contact-rich manipulation [79], while ScrewMimic models bimanual tasks as constrained screw motions for learning from human videos [3].
-However, tactile methods require extra hardware, and the screw motion assumption may not hold.
-Another line of work by Wen et al.
-[70] aims to learn category-level representations from a single demo to transfer pose trajectories across similar objects [70], but assumes a static target (e.g., battery slot), limiting real-world applicability in real-world tasks where the spatial configuration of objects may change.
-Unlike these methods, our method only requires natural human data, without assuming access to tactile sensors or requiring constrained screw motion models, making it significantly more cost-effective, scalable, and accessible.
-II-B Cross-Embodiment Policy Learning
-Cross-embodiment policy learning enables robots to transfer policies across embodiments, such as those with different kinematic structures [25, 52, 77].
-Prior work has explored conditioning policies on embodiment representations using multi-embodiment datasets [18, 69, 80, 13, 24, 52, 2], but these approaches often face challenges in effectively leveraging human interaction data.
-Recent approaches utilize human data, such as estimating point flow from human video [72] or generating latent plan or high-level plans from human data [37, 68].
-However, their dependence on robot data for low-level control limits scalability when robot data is expensive or difficult to collect.
-Additionally, prior works highlight the importance of masking human and robot embodiments for visual consistency [4, 31].
-However, Bahl et al.
-[4] relies on predefined motion primitives, while Kareer et al.
-[31] still requires robot data with human data as augmentation.
-Our approach adopts a similar masking idea but enables robots to learn freely, even agile motions from human videos, without any robot data.
-Another line of work tokenizes observation inputs and action outputs into a unified transformer network, enabling generalized policy learning across embodiments [59, 76].
-However, these methods require large models and extensive datasets, making them resource-intensive and time-consuming to train, and lack the capability for direct policy transfer between embodiments.
-Overall, reliance on robot-specific hardware and data restricts scalability and accessibility.
-In contrast, our method leverages natural human data, eliminating the need for robot data as a training source.
-II-C Cross-Viewpoint Policy Learning
-Viewpoints can change as robots manipulate or interact with their environment.
-Existing approaches address viewpoint changes by learning view-invariant latent representations [58, 11], leveraging equivariant 3D representations [63, 23, 89], or augmenting input data with diverse viewpoints [51, 87].
-An early attempt by Sharma et al.
-[61] introduced a hierarchical model that generates first-person sub-goals from third-person demonstrations via a high-level module, while a low-level embodiment-specific module predicts actions.
-Recent work by Chen et al.
-[12] leveraged the SAM model [32] and ControlNet [83] to transform the viewpoint of one robot into another, performing view synthesis with ZeroNVS [54] to augment datasets and enhance policy learning across robots and viewpoints.
-Another approach by Yuan et al.
-[82] combined reinforcement learning with multi-view representation learning and a Spatial Transformer Network (STN) to improve policy robustness in visually complex environments.
-However, these methods face limitations such as scalability issues in hierarchical models, the complexity of controlled image editing, high computational costs in view synthesis, and the need to build sophisticated simulators for reinforcement learning when adapting to diverse environments.
-Compared to previous work, our method enables efficient and scalable cross-viewpoint transfer from human data by leveraging two RGB images with the MASt3R [34] foundation model for 3D reconstruction and employing Gaussian splatting for fast and cost-effective novel view synthesis.
-III Problem Statement
-We formulate the robotic manipulation task as a Markov Decision Process, where the goal is to learn a policy $\pi:\mathcal{O}^{r}\to\mathcal{A}$ that enables a robot to perform a given task.
-The robot’s observation space, $\mathcal{O}^{r}$, consists of a single-view RGB image $I^{r}\in\mathbb{R}^{128\times 128\times 3}$ and proprioception data $x^{r}\in SE(3)$.
-To train the policy, we use an imitation dataset composed of $N$ human play, denoted as $D=\{(O_{0}^{h},O_{1}^{h},\dots)\}_{n=1}^{N}$.
-Each human play observation $\mathcal{O}^{h}$ consists of two RGB images captured from different viewpoints: $\mathcal{O}^{h}=\{I^{v1},I^{v2}\}$, where $I^{vi}\in\mathbb{R}^{480\times 640\times 3}$.
-We preprocess the dataset to get the action by using a 6D pose estimation and tracking model, resulting in $D=\{(O_{0}^{h},a_{0},O_{1}^{h},a_{1},\dots)\}_{n=1}^{N}$, where each action $a\in SE(3)$.
-Our objective is to train a robot to perform the same task as demonstrated in the human dataset while overcoming the embodiment differences.
-We assume that the tool is rigid, the tool and the end effector are rigidly attached during deployment, and the transformation between the tool and the end effector is estimated once before deployment.
-Our approach aims to allows the robot generalize task execution across embodiments while retaining the core skills demonstrated by humans.
-IV Our Robot Policy Learning Framework
-Our framework enables the direct transfer of human play data into deployable robot policies. It is designed to fulfill the following key objectives:
-  • Support for Dynamic and High-Precision Tasks: Human play, with its inherent fluidity, enables the execution of highly dynamic tasks.
+Human manipulation data was collected using two cameras and processed through the foundation model MASt3R [61] to generate 3D reconstructions.
+Using 3D Gaussian splatting, we sampled novel views to augment the dataset.
+Human hands were segmented to create embodiment-agnostic observations as policy inputs.
+For action labeling, FoundationPose [62] estimated the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$, which was transformed into task space, $T_{\text{tool}}^{\text{task}}$.
+A diffusion model was then trained as the visuomotor policy.
+Problem Statement: We formulate robotic manipulation as a Markov Decision Process (MDP), where the goal is to learn a policy $\pi:\mathcal{O}^{r}\to\mathcal{A}$ that enables a robot to perform a given task.
+The robot’s observation space $\mathcal{O}^{r}$ consists of a single-view RGB image $I^{r}\in\mathcal{I}^{r}$ and proprioceptive data $x^{r}\in SE(3)$, where each $I^{r}$ is a tensor in $\mathbb{R}^{128\times 128\times 3}$.
+We train the policy using an imitation dataset of $N$ human demonstrations, $D={(\mathcal{O}_{0}^{h},\mathcal{O}_{1}^{h},\dots)}_{n=1}^{N}$, where each $\mathcal{O}^{h}=\{I_{v1}^{h},I_{v2}^{h}\}$ contains two RGB images captured from different viewpoints and each $I_{vi}^{h}\in\mathcal{I}^{h}$ is a tensor in $\mathbb{R}^{480\times 640\times 3}$.
+We preprocess the dataset to infer actions using a 6D pose estimation and tracking model, resulting in $D=\{(\mathcal{O}_{0}^{h},a_{0},\mathcal{O}_{1}^{h},a_{1},\dots)\}_{n=1}^{N}$, where each action $a\in SE(3)$.
+To bridge the embodiment gap between humans and robots, we assume the tool is rigidly attached to both the human hand (implicitly) and the robot end-effector (explicitly), with a fixed transformation estimated prior to deployment.
+Under this setup, the robot can reproduce human-demonstrated tool trajectories, enabling policy transfer across embodiments while preserving task-relevant behaviors (Figure 2).
+Tool-Centric Demonstrations for Robot Manipulation: We leverage the fact that both humans and robots can operate the same physical tools to facilitate policy learning.
+Tools serve as a shared interface for interacting with objects, enabling the direct transfer of human demonstrations with minimal embodiment-specific adaptation.
+Unlike prior work focused on grasping or pick-and-place tasks [14, 13, 63], our approach enables robots to perform complex interactions using everyday tools.
+Our formulation abstracts actions to the tool pose, reducing morphological dependence and promoting policy generalization across embodiments.
+It also simplifies data collection by eliminating the need for robot-specific demonstrations.
+Humans can naturally manipulate tools by hand without extra instrumentation.
+For deployment, robots either rigidly grasp the tool, as demonstrated with a Kinova Gen3 arm, or attach it using a custom fast tool changer described in Appendix B.2.1 and shown in Figure 10, compatible with ISO 9409-1-50-4-M6 flanges.
+Perception Alignment Across Embodiments: To enable cross-embodiment policy transfer, we align human and robot observations within a shared visual space $\mathcal{I}^{s}$ by applying a feature extraction function $g:\mathcal{I}^{h}\cup\mathcal{I}^{r}\to\mathcal{I}^{s}$.
+We instantiate $g$ with Grounded-SAM [64], using prompts such as “human hand” and “robot arm” to mask out embodiment-specific regions—human hands during training and robotic arms during deployment.
+Masking out these regions ensures that only task-relevant visual information (e.g., tools and objects) remains visible in both phases.
+By minimizing visual discrepancies between human and robot data, the feature extraction process reduces embodiment-specific bias and improves generalization across embodiments.
+3D-Aware View Augmentation: We use cameras for data collection due to their widespread availability—over 7.14 billion smartphones are equipped with them [65].
+However, single-camera setups suffer from scale ambiguity and limited 3D perception and are sensitive to viewpoint changes.
+3D Reconstruction: To address this, we use MASt3R [61], an image-matching model that reconstructs accurate 3D environments from just two RGB images—eliminating the need for depth sensors, which are less common and more power-hungry.
+Two cameras suffice to avoid scale ambiguity inherent in monocular settings.
+MASt3R produces high-quality point clouds without requiring known camera extrinsics or intrinsics by globally aligning multi-view features.
+View Synthesis and Augmentation: 3D Gaussian splatting synthesizes novel viewpoints from the reconstructed scene, allowing the robot to observe interactions from multiple angles—even when only two views are available.
+The resulting perspectives augment the training data, increasing visual diversity and improving policy learning.
+To further enhance robustness and generalization, random cropping is applied, following diffusion policy [25, 26].
+Tool-Centric Action Representation and Policy Deployment: To support general tool usage, we propose a task-frame, tool-centric action representation denoted as $T_{\text{tool}}^{\text{task}}$, which describes the tool’s motion independently of human or robot morphology, camera pose, or base configuration.
+This invariant formulation enables robust policy transfer across different embodiments and viewpoints.
+As shown in Figure 3, the tool’s pose is first estimated in the camera frame using a 6D pose estimation model (e.g., FoundationPose [62]) as $T_{\text{tool}}^{\text{camera}}$, and then transformed into the task frame:
+$T_{\text{tool}}^{\text{task}}=T_{\text{camera}}^{\text{task}}T_{\text{tool}}^{\text{camera}},$
+where $T_{\text{camera}}^{\text{task}}$ denotes the transformation from the camera to the task frame.
+Figure 3: Coordinate System Diagram. The diagram shows the Camera, Tool, Task space, End-Effector (EEF), and Base frames. The action is represented as $T_{\text{tool}}^{\text{task}}$.
+A diffusion policy [25] maps a single-view RGB image to a predicted SE(3) action $T_{\text{tool}}^{\text{task}}$.
+At deployment, the robot command is computed by converting the prediction to the end-effector frame.
+For stationary robots, the task frame aligns with the base frame; for mobile platforms, base movement is compensated using $T_{\text{task}}^{\text{base}}$.
+The resulting end-effector pose is given by:
+$T_{\text{eef}}^{\text{base}}=T_{\text{task}}^{\text{base}}T_{\text{tool}}^{\text{task}}T_{\text{eef}}^{\text{tool}},$
+where $T_{\text{eef}}^{\text{tool}}$ is the known fixed transform between the tool and the robot end-effector.
+4 Policy Evaluations
+Our evaluations aim to assess our framework across three dimensions: reliability (how consistently and successfully the learned policies perform), execution efficiency (how smooth and natural the resulting behaviors are), and versatility (how well the framework adapts to diverse tasks and generalizes across conditions).
+Experimental Tasks Overview: We evaluate five real-world robotic tasks on Kinova Gen3 and UR5e robots, involving precision manipulation, dynamic object handling, and dexterous tool use.
+Policies use RGB inputs from RealSense D415 cameras and handle variations in object positions and camera poses.
+Tasks include: (1) Nail Hammering – Precise striking of a small target, (2) Meatball Scooping – Contact-sensitive rolling object manipulation, (3) Pan Flipping – Fast, dynamic flipping with varied objects, (4) Wine Balancing – Gravity-aware placement into an unstable rack, and (5) Soccer Ball Kicking – Dynamic interception and obstacle avoidance.
+Full details in Appendix B.1.
+Baselines: We evaluate the effectiveness and efficiency of learning directly from human manipulation videos without relying on robot-generated data.
+We benchmark against a diffusion policy trained on robot demonstrations and UMI [14], a hand-held gripper method.
+Robot demonstrations are collected using SpaceMouse or Gello [19] under identical time budgets.
+Additionally, we conduct ablations to assess random cropping before policy training, novel view synthesis data augmentation, and embodiment segmentation.
+To further illustrate the advantages of our approach, we compare trajectory rollouts for a meatball-scooping episode, highlighting how our method is more sample-efficient and less prone to distribution shifts by eliminating excessive waypoints.
+Evaluation Metrics: During testing, we introduce two types of variations: (1) randomizing the initial spatial configurations of objects in each task to assess policy generalization, and (2) varying camera positions to evaluate the robustness of policies to different viewpoints.
+All methods, including the baseline and ablation variants, are tested under the same conditions.
+Performance is evaluated using two metrics: success rate, which measures the proportion of successfully completed task trials and reflects policy effectiveness, and task completion time, which captures the average duration to complete tasks and reflects policy efficiency.
+5 Experiment Results
+Capabilities and Effectiveness: Table 1 summarizes our real-world results, showing that our framework consistently achieves higher success rates across all tasks compared to baselines.
+We also compare against the stronger hand-held gripper baseline UMI [14] (Table 2).
+In our default setup, SLAM-based mapping failed due to low environmental texture, so we added a textured background to support reliable mapping for UMI.
+For the nail hammering task, we evaluated UMI with 25 demonstrations (matching our collection time) and 100 demonstrations (to assess ideal performance).
+UMI fails all 13 trials with 25 demonstrations but succeeded with 100.
+It was also inapplicable
+Table 1: Task Success Rates and Completion Times. Success rates show completed trials over total. “DP” is trained on teleoperation; “Not Feasible” means failure.
+Task Method Success Time (s)
+Hammer DP 0/13
+Ours 13/13 11.0
+Scoop DP 5/12 42.0
+Ours 10/12 12.4
+Pan: Egg DP Not Feasible
+Ours 12/12 1.5
+Pan: Bun DP Not Feasible
+Ours 9/12 1.9
+Pan: Patty DP Not Feasible
+Ours 10/12 2.3
+Wine Balance DP Not Feasible
+Ours 8/10 30.9
+Soccer Kick DP Not Feasible
+Ours 6/10 2.0
+to wine balancing and pan flipping due to contact and inertial challenges, and struggled in soccer kicking due to localization failures.
+In contrast, our method demonstrates reliable performance across all tasks: accurately detecting spatial locations (nail hammering, meatball scooping), performing high-speed motions (pan flipping), precisely inserting wine bottles, and swiftly reacting in soccer kicking.
+This strong performance is enabled by collecting significantly larger and more diverse episodes within the same data collection timeframe (see Section 6), enabling robust policy training.
+Our approach overcomes limitations of teleoperation tools like Gello and SpaceMouse, enabling data collection for scenarios they struggle to handle.
+Qualitative policy rollouts are shown in Figure 9 in Appendix.
+Policy Execution Efficiency: Our framework demonstrates high execution efficiency, achieving shorter task completion times and smoother action trajectories than baselines (Table 1).
+This efficiency stems from the natural human data, which captures the fluidity and speed of real-world activities, resulting in higher-quality training trajectories.
+In contrast, teleoperated demonstrations
+Table 2: Task success rates comparing our method with the hand-held gripper-based method on Nail Hammering.
+Method Demo Time & Count Success
+UMI [14] 180s (25) 0/13
+UMI 720s (100) 13/13
+Ours 180s (40) 13/13
+often produce slower, less natural motions, limiting dynamic performance.
+By leveraging more realistic data, our framework accelerates execution while enhancing motion quality, making it suited for real-world applications.
+See Appendix C for a detailed comparison of policy rollouts.
+Benefits of Tool-Based Action Representation in Task Space: Using the tool pose in the camera frame works with a static camera but fails under camera movement due to unreliable real-time tracking and incorrect end-effector positioning in the base frame.
+Similarly, representing actions in the base frame fails under base movement due to the assumption of a fixed base-to-workspace transform.
+In contrast, representing actions in task space is invariant to both camera and base movement, enabling robust execution even under large viewpoint shifts and base movements.
+Figure 4: Policy Testing Across Camera Poses in Nail Hammering.
+(a) Camera poses for data collection and evaluation.
+(b-d) Performance ranges for methods trained with/without random cropping (RC) and view augmentation (VA).
+Figure 5: Robustness to Camera and Base Movement.
+(a) Camera Pose Robustness: The policy demonstrated the ability to handle camera shaking across three tasks—meatball scooping, nail hammering, and pan flipping.
+The first row shows the camera view, while the second row provides a scene overview with the shaking motion.
+(b) Robot Base Robustness: The policy successfully compensated for base shaking, even when the shaking frequency exceeded the robot’s control frequency.
+(c) Chicken Head Stabilization: At lower base movement frequencies, the end effector displayed a stabilization effect similar to a chicken’s steady head.
+(d) Combined Robustness: The policy maintained task performance under simultaneous camera and base shaking.
+Effects of Random Cropping and View Augmentation: Our experiments show that random cropping (RC) and view augmentation (VA) together enhance policy robustness to camera pose variations.
+RC improves resilience to minor perturbations such as small movements or shaking, while VA exposes the model to a broader distribution of viewpoints during training.
+We evaluated these techniques on the nail hammering task (Figure 4), comparing three models: one trained with both RC and VA, one with RC only, and one without either.
+The combined use of RC and VA significantly expands the range of camera configurations under which the policy can successfully operate.
+Generalization:
+Spatial Generalization: We evaluated spatial generalization by varying initial conditions across tasks: nail positions for hammering, meatball locations for scooping, goalkeeper setups for soccer ball kicking, and object poses across the pan for flipping (illustrated in Figure 11 in Appendix).
+Object Generalization: Our method generalizes effectively to different objects in the pan-flipping task, including the egg and burger bun seen during training, and a 3D-printed meat patty (illustrated in Figure 11, second column, in Appendix).
+The policy learns to tilt the pan to slide the object into a corner, then flick it to achieve a successful flip, enabling robust generalization across object types.
+Tool Generalization: We evaluated tool generalization by testing the policy with five different pans: large, medium, small, tiny, and square.
+The policy was trained using demonstrations with the large, medium, and square pans and evaluated on all five, with 12 trials per pan under varying initial configurations (illustrated in Figure 13 in Appendix).
+It achieved high success rates on the trained pans (large and medium).
+Performance declined on smaller pans due to limited surface area, and on the square pan due to shallow edges causing the bun to slide out during flipping.
+Figure 6: Human Perturbation Robustness.
+The robot handles human-induced perturbations across three tasks: (1) In nail hammering, it tracked a manually moved nail; (2) In meatball scooping, it located and scooped new meatballs thrown mid-task; and (3) In egg flipping, it recovered the egg after human repositioning.
+Robustness:
+Camera Pose Robustness:
+We evaluated the policy’s ability to handle camera pose variations by introducing camera shaking in three tasks: meatball scooping, nail hammering, and pan flipping (Figure 5(a)).
+The first row shows the camera view, and the second shows the scene overview and shaking motion.
+Despite disturbances, the policy consistently completed all tasks, enabled by random cropping during training, improving adaptation to partial views and minor visual changes.
+Robot Base Robustness: To assess robustness to base movement, we manually shook the robot base during execution (Figure 5(b)).
+When the shaking frequency exceeded the control frequency, the end effector oscillated with the base; however, the task-space action design enabled compensation and successful task completion.
+As shown in Figure 5(d), the policy also maintained effectiveness under simultaneous camera and base shaking.
+Chicken Head Stabilization: At lower shaking frequencies, where the perturbation was slower than the robot’s control loop, the end effector exhibited a stabilization behavior similar to a chicken’s head [66] (Figure 5(c)), maintaining steady control during mild base movements.
+Human Perturbation Robustness: We evaluated resilience to human interventions (Figure 6).
+The robot tracked moving nails, adapted to new meatballs thrown in mid-task, and re-flipped repositioned eggs, demonstrating robustness to real-time disturbances.
+Figure 7: Data Collection Efficiency and Reliability.
+(a) Human hands excel in manipulation tasks, leveraging natural and intuitive efficiency.
+(b) Failure cases for Gello and Spacemouse include insufficient speed, lack of tactile feedback during data collection, safety stops, collisions, teleoperation delays, and difficulty handling high-speed or complex tasks.
+(c) Failure cases for handheld grippers such as UMI [14], where issues arise from tool slippage due to inertia or displacement caused by contact forces.
+6 Data Collection Efficiency and Affordability
+We compare data collection methods for robot imitation learning across throughput, reliability, cost, usability, and precision. Full quantitative and qualitative analysis is provided in Appendix D.
+Data Collection Efficiency: Leveraging the natural dexterity and intuitive control of human hands (Figure 7(a)), our method achieves significantly higher data collection throughput compared to traditional methods.
+Quantitatively, human demonstrations reduce data collection time by 73% for nail hammering and 81% for meatball scooping relative to teleoperation methods (Gello, Spacemouse), as shown in Figure 8.
+Compared to handheld grippers (e.g., UMI [14]), human demonstrations are 41% faster in nail hammering, and further succeed in dynamic tasks where both teleoperation and handheld methods consistently fail.
+This efficiency enables high-throughput, low-variance data collection critical for scalable robot learning.
+Reliability: In contrast, Figures 7(b) and (c) highlight typical failures for Gello, Spacemouse, and handheld grippers, including collisions, tool slippage, safety stops, and failures in dynamic or high-precision tasks.
+Teleoperation tools frequently suffer from latency, lack of tactile feedback, and difficulty handling rapid motions, leading to inconsistent trajectories and poor-quality demonstrations.
+Handheld grippers are prone to inertial slippage and loss of tool control under high forces, further limiting their applicability.
+Figure 8: Quantitative Comparison of Data Collection Methods.
+Human hands reduce data collection time by 73% for nail hammering and 81% for meatball scooping, while maintaining low variation.
+Teleoperation fails in dynamic and high-precision tasks.
+In nail hammering, human hands are 41% faster than UMI [14], which also struggles with dynamic and low-texture environments.
+Cost and Usability: Our method requires no specialized hardware, incurring zero additional cost compared to teleoperation devices, handheld grippers, and VR/AR equipment, all of which demand substantial investment and maintenance.
+A detailed discussion is provided in Appendix D, with a direct comparison summarized in Table 4.
+The intuitive nature of human demonstrations further lowers the expertise barrier, making large-scale, high-quality data collection accessible without costly infrastructure.
+By eliminating hardware dependence, minimizing operational failures, and dramatically improving throughput and precision, our human-centric data collection framework enables scalable, efficient, and reliable robot imitation learning.
+7 Conclusion
+In this work, we presented a framework that enables robust policy training for diverse tool-use tasks by learning from human manipulation videos and bridging the embodiment gap they introduce.
+Unlike traditional data collection methods, which are costly and hardware-dependent, our approach democratizes data collection by eliminating the need for specialized equipment or technical expertise, making large-scale robot learning accessible and scalable.
+We validated the framework across challenging tasks, including nail hammering, meatball scooping, pan flipping, wine bottle balancing, and soccer ball kicking, demonstrating superior performance, robustness to variations in camera poses and base movements, and adaptability across 6-DOF and 7-DOF robots.
+By improving accessibility, scalability, and reliability, our work lays a strong foundation for advancing robotic manipulation in complex, real-world scenarios.
+8 Limitations and Future Work
+Our framework has certain limitations.
+First, the perception pipeline relies on FoundationPose for extracting the tool’s pose during manipulation.
+Errors in pose estimation may occasionally require data recollection, adding time and effort.
+Improving the reliability of the perception pipeline through more robust pose estimation algorithms or self-correction mechanisms is a promising direction for future work.
+Second, for novel view augmentation, significant noise and reduced realism are observed when augmented views deviate too far from the collected camera views, which can hinder policy performance.
+Future efforts could focus on leveraging advanced rendering techniques to enhance the realism of augmented views and improve policy generalization.
+Third, we assume the tool is rigidly attached to the robot’s end effector; however, in real-world, contact-rich manipulation, minor shifts may occur, potentially affecting performance.
+Addressing this issue by incorporating tactile sensing could improve performance in contact-intensive tasks.
+Additionally, our method assumes a rigid tool and does not account for flexible or soft tools.
+Future work could explore using flexible representations for tool state estimation to better handle deformable tools in real-world manipulation scenarios.
+Acknowledgments
+We thank Professor Wenzhen Yuan, Xiaoyu Zhang, Yucheng Mo, Yilong Niu, Hyoungju Lim, and Amin Mirzaee for their support and assistance in reproducing the UMI.
+We thank Professor Saurabh Gupta, Professor Junyi Geng, Lujie Yang, Neeloy Chakraborty, Hongkai Dai, Jacob Wagner, Shaoxiong Yao, Wei-Cheng Huang for their insightful feedback and suggestions.
+This work was supported by ZJU-UIUC Joint Research Center Project No.
+DREMES 202003, funded by Zhejiang University.
+This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
+Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
+Additionally, this work used NCSA Delta GPU at NCSA through allocation CIS240753 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S.
+National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
+The robot experiments were carried out in part in the Center for Autonomy Robotics Laboratories at the University of Illinois Urbana-Champaign.
+APPENDIX
+Appendix A Design Choice
+A.1 Key Capabilities and Practical Benefits
+Our framework enables the direct transfer of human manipulation data into deployable robot policies. It is designed to fulfill the following key objectives:
+  • Support for Dynamic and High-Precision Tasks: Human manipulation, with its inherent fluidity, enables the execution of highly dynamic tasks.
 Examples include flipping an egg in a pan or performing other actions that require swift, accurate, and natural motions — challenges that are often difficult to address with traditional teleoperation systems or handheld grippers.
@@
 Our approach removes technical barriers associated with 3D printing and other hardware setups, fostering a seamless, user-friendly experience for data collection.
-The following sections outline the framework’s design (Figure 2), its underlying principles, and logistical considerations for development and deployment.
-IV-A Tool Usage for Data Collection and Manipulation
-Humans naturally and intuitively use tools for everyday tasks such as cooking, eating, cleaning, and interacting with the world.
-Tools act as extensions of human actions, enabling diverse interaction with objects.
-The natural relationship between tools and objects provides an ideal interface for training robots to mimic human actions using tools, with minimal gap between human and robot tool usage.
-While grasping and pick-and-place tasks have been extensively studied in previous works, our work focuses on enabling robots to use the same tools humans commonly employ to interact with their environment effectively with the following benefits:
-  • Minimized Embodiment Gap: Abstracting actions to the tool pose reduces morphological dependency, enabling policies to generalize across embodiments.
-  • Scalable Data Collection: Simplifies the data collection process by eliminating the need for costly robot-specific demonstrations, making our method more accessible.
-During data collection, humans can naturally use tools with their hands without requiring additional devices. For robot deployment, the tool can be attached to the robot in two ways:
-  • Rigid Grasping: Grasping or picking, which has been extensively studied in prior works, is demonstrated in our Kinova Gen3 robot experiments and involves the robot securely grasping the tool.
-  • Customized Fast Tool Changer: Designed for versatility, the tool changer is compatible with any robot that uses the ISO 9409-1-50-4-M6 flange, as demonstrated in our UR5e experiments.
-IV-B Embodiment-Agnostic Perception
-To encourage cross-embodiment transfer, we adopt a strategy that reduces the perception gap between the training and deployment phases.
-During training, human play data is collected, featuring human hands interacting with tools and objects.
-In deployment, robots execute the learned tasks.
-As showcased in our experiments, the visual differences between human hands and robotic end-effectors can introduce discrepancies that hinder generalization.
-To address this, we employ Grounded-SAM [49] to segment and mask out the embodiments in each phase.
-During training, human hands are masked, while during deployment, the robotic embodiments are masked, which ensures that the remaining parts of the scene in both training and testing phases appear visually similar.
-By aligning perception across embodiments, the framework mitigates distractions caused by embodiment-specific features, enabling better generalization to human-to-robot policy transfer.
-IV-C View Augmentation
-We use cameras for data collection due to their availability.
-With approximately 7.14 billion smartphones equipped with cameras, our approach can scale effectively [30].
-However, using data from a single camera introduces challenges such as a lack of 3D perception and sensitivity to camera pose.
-3D Reconstruction
-To address these issues, we use MASt3R [34], an image-matching model that reconstructs accurate 3D environments from two RGB images.
-This eliminates the need for additional depth sensors, which are less common and consume more power compared to RGB sensors.
-We use two cameras, which is the minimum requirement for 3D reconstruction, to capture demonstration data.
-Using only one camera can lead to scale ambiguity in monocular settings, where poses can be scaled by an arbitrary, scene-dependent factor.
-By capturing images from two viewpoints, MASt3R reconstructs a 3D point cloud without requiring camera extrinsics or intrinsics, and globally aligns point maps within a multi-view 3D reconstruction framework.
-The process results in high-quality 3D representations.
-Data Augmentation
-Using 3D Gaussian splatting, we model the scene and synthesize novel viewpoints from human interaction data, which effectively augments the dataset, generating additional perspectives even if the training data was captured from only two views.
-These synthetic viewpoints provide the robot with a multi-angle understanding of the scene, allowing the policy to be trained on a more diverse and comprehensive set of visual inputs.
-Additionally, we apply random cropping to the images for data augmentation before feeding them into the policy network, following the approach from diffusion policy [14, 15].
-Random cropping further improves the method’s robustness, enabling the policy to generalize better to variations in visual inputs.
-IV-D Action Representation for Tool Manipulation
-To support general tool usage, we propose a task-frame, tool-centric action representation denoted as $T_{\text{tool}}^{\text{task}}$.
-This representation focuses on the tool being manipulated, independent of human or robot morphology or camera pose.
-A visual depiction of the coordinate systems is provided in Figure 3.
-Using a 6D pose estimation model (e.g., FoundationPose [71]), we determine the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$.
-To make the policy robust to camera and base movement, we transform this into the task frame:
-$T_{\text{tool}}^{\text{task}}=T_{\text{camera}}^{\text{task}}T_{\text{tool}}^{\text{camera}},$
-where $T_{\text{camera}}^{\text{task}}$ represents the transformation from the camera to the task frame.
-TABLE I: Benchmark Attributes of Real-World Tasks.
+Appendix B Detailed Experiment Setup
+B.1 Task Descriptions
+Nail Hammering: The task involves hammering a 3D-printed nail, requiring the robot to locate the nail, draw back the hammer, and strike the nail tip accurately.
+With a diameter of less than 15.5 mm, the nail tip demands high precision.
+Challenges include localizing the nail tip precisely and planning effective hammer trajectories.
+To evaluate generalization, the initial position of the nail is varied across different spatial configurations.
+We collected 180 seconds of data (40 episodes) from a single participant.
+Meatball Scooping: In this task, the robot must use a spoon to scoop a meatball from a pan and transfer it to a bowl.
+This task is challenging due to the complex dynamics of the meatball, which can roll unpredictably within the pan.
+Additionally, the interaction between the spoon and the meatball requires careful control, as improper contact can cause the meatball to slip or escape the spoon.
+We randomize the initial position of the meatball within the pan to test its generalization capability.
+We collected 340 seconds of data (50 episodes) from a single participant.
+Pan Flipping (Egg, Burger Bun, Meat Patty): The objective of this task is to use a pan to flip various objects, such as an egg, a burger bun, and a meat patty.
+The task is challenging due to its high-speed dynamics, requiring the robot to overcome gravity and accurately manage the interaction between the pan and the objects.
+Each object differs in weight, shape, and texture, adding further complexity.
+This task evaluates the policy’s ability to handle fast, contact-rich interactions and adapt to diverse object types.
+To increase variability, the initial positions of the objects within the pan are randomized.
+Furthermore, the rapid and dynamic nature of the task makes it unsuitable for classical demonstration collection methods, highlighting the advantages of using bare-handed human videos for data collection.
+We collected 50 seconds of data (38 episodes) from a single participant using three different pans and two object types.
+Wine Balancing: In this task, the robot needs to use a hook to lift a wine bottle and carefully insert it into an unstable, zero-gravity wine rack.
+The task is challenging due to the precise control required to suspend the bottle in mid-air and counteract gravitational forces effectively.
+Any over-insertion or under-insertion will cause the bottle to lose balance.
+To constrain the horizontal movement of the rack, screws were added as obstacles to limit lateral motion.
+No additional variability was introduced.
+We collected 223 seconds of data (15 episodes) from a single participant.
+Soccer Ball Kicking: In this task, the robot must use a golf club to kick a ball that slides into a field and direct it into the goal.
+To increase the challenge, a 3D-printed row of players serves as obstacles between the robot and the goal.
+The task is difficult because the robot must accurately intercept the moving ball, strike it with the correct force and direction, and ensure it avoids obstacles before reaching the goal.
+The position of the player obstacle varies.
+We collected 78 seconds of data (20 episodes) from a single participant.
+Table 3: Benchmark Attributes of Real-World Tasks.
 These benchmarks evaluate the precision, adaptability, and capability of our framework to address tasks requiring high precision, handling extreme dynamics, utilizing extrinsic dexterity, performing in contact-rich scenarios, and overcoming gravity.
@@
 Task 5: Soccer Ball Kicking
-IV-E Training and Deploying Robot Policies
-We use diffusion policy [14] as our policy representation to predict $T_{\text{tool}}^{\text{task}}$, trained using the ACCESS system [7].
-During deployment, for stationary robots, the task frame aligns with the base frame.
-For robots with moving base, base movement is compensated using $T_{\text{task}}^{\text{base}}$.
-The final end-effector pose in the robot base frame, used as the control command for the robot controller, is computed as:
-$T_{\text{eef}}^{\text{base}}=T_{\text{task}}^{\text{base}}T_{\text{tool}}^{\text{task}}T_{\text{eef}}^{\text{tool}},$
-where $T_{\text{eef}}^{\text{tool}}$ is the fixed transformation between the tool and the robot end-effector.
-Figure 3: Coordinate System Diagram.
-This diagram illustrates the Camera frame, Tool frame, Task space frame, End-Effector (EEF) frame, and Base frame.
-We use $T_{\text{tool}}^{\text{task}}$ as the action representation for the visuomotor policy output.
-V Policy Evaluations
-Our experimental evaluations aim to assess the effectiveness of our framework for deploying robot policy learning across three key dimensions:
-  • Capabilities: What skills can our framework enable robots to acquire, and how robust are the policies under challenging conditions like a moving camera or base?
-  • Effectiveness: How well does our framework achieve its objectives outlined in Section IV?
-Can it support reliable and scalable policy learning for complex tasks while streamlining data collection for highly dynamic, contact-rich, or dexterous scenarios?
-  • Policy Execution Efficiency: How efficiently do the trained policies execute tasks?
-Does our framework enable smoother and more natural motion trajectories?
-Can it achieve faster and more fluid task completion compared to baseline methods?
-To evaluate our framework for learning from direct human play, we developed a set of real-world robotic task domains using two embodiments: Kinova Gen3 and UR5e.
-These tasks are designed to test various aspects of policy capabilities, efficiency and effectiveness.
-Table I summarizes the key characteristics of these tasks.
-Visual input for the policies is provided by two RealSense D415 cameras.
-We describe the tasks in detail, highlighting their challenges and the specific capabilities we aim to test.
 Figure 9: Policy Rollouts.
 We evaluate diverse real-world tasks: nail hammering (precision in locating a nail tip), meatball scooping (slippery object, constrained environments), pan flipping (extremely dynamic, high-speed, contact-rich), wine balancing (precise control of unstable objects), and soccer ball kicking (dynamic object handling, goal-directed actions).
-Nail Hammering: The task involves hammering a 3D-printed nail, requiring the robot to locate the nail, draw back the hammer, and strike the nail tip accurately.
-With a diameter of less than 15.5 mm, the nail tip demands high precision.
-Challenges include localizing the nail tip precisely and planning effective hammer trajectories.
-To evaluate generalization, the initial position of the nail is varied across different spatial configurations.
-We collected 180 seconds of data (40 human play episodes) from a single participant.
-Meatball Scooping: In this task, the robot must use a spoon to scoop a meatball from a pan and transfer it to a bowl.
-This task is challenging due to the complex dynamics of the meatball, which can roll unpredictably within the pan.
-Additionally, the interaction between the spoon and the meatball requires careful control, as improper contact can cause the meatball to slip or escape the spoon.
-We randomize the initial position of the meatball within the pan to test its generalization capability.
-We collected 340 seconds of data (50 human play episodes) from a single participant.
-Pan Flipping (Egg, Burger Bun, Meat Patty): The objective of this task is to use a pan to flip various objects, such as an egg, a burger bun, and a meat patty.
-The task is challenging due to its high-speed dynamics, requiring the robot to overcome gravity and accurately manage the interaction between the pan and the objects.
-Each object differs in weight, shape, and texture, adding further complexity.
-This task evaluates the policy’s ability to handle fast, contact-rich interactions and adapt to diverse object types.
-To increase variability, the initial positions of the objects within the pan are randomized.
-Furthermore, the rapid and dynamic nature of the task makes it unsuitable for classical demonstration collection methods, highlighting the advantages of using bare-handed human play for data collection.
-We collected 50 seconds of data (38 human play episodes) from a single participant using three different pans and two object types.
-Wine Balancing: In this task, the robot needs to use a hook to lift a wine bottle and carefully insert it into an unstable, zero-gravity wine rack.
-The task is challenging due to the precise control required to suspend the bottle in mid-air and counteract gravitational forces effectively.
-Any over-insertion or under-insertion will cause the bottle to lose balance.
-To constrain the horizontal movement of the rack, screws were added as obstacles to limit lateral motion.
-No additional variability was introduced.
-We collected 223 seconds of data (15 human play episodes) from a single participant.
-Soccer Ball Kicking: In this task, the robot must use a golf club to kick a ball that slides into a field and direct it into the goal.
-To increase the challenge, a 3D-printed row of players serves as obstacles between the robot and the goal.
-The task is difficult because the robot must accurately intercept the moving ball, strike it with the correct force and direction, and ensure it avoids obstacles before reaching the goal.
-The position of the player obstacle varies.
-We collected 78 seconds of data (20 human play episodes) from a single participant.
-Baselines: The primary focus of this work is to demonstrate the effectiveness and efficiency of learning directly from human play without relying on robot-generated data.
-We compare our approach against two baselines: a diffusion policy trained on robot demonstrations, and UMI [16], a hand-held gripper-based data collection method.
-The robot demonstration dataset is collected using either a SpaceMouse or Gello interface under the same data collection time.
-Additionally, we perform ablation studies to analyze the impact of key components, such as random cropping of images before policy training, novel view synthesis-based data augmentation, and the effects of embodiment segmentation.
-To further illustrate the advantages of our approach, we compare trajectory rollouts for a meatball-scooping episode, highlighting how our method is more sample-efficient and less prone to distribution shifts by eliminating excessive waypoints.
-Evaluation Metrics: During testing, we introduce two types of variations: (1) randomizing the initial spatial configurations of objects in each task to assess policy generalization, and (2) varying camera positions to evaluate the robustness of policies to different viewpoints.
-All methods, including the baseline and ablation variants, are tested under the same conditions.
-Performance is evaluated using two metrics: success rate, which measures the proportion of successfully completed task trials and reflects policy effectiveness, and task completion time, which captures the average duration to complete tasks and reflects policy efficiency.
-TABLE II: Task Success Rates and Completion Times.
-Success rates are the number of successful trials out of total episodes, and average completion times are based on successful trials.
-“DP” refers to the diffusion policy trained on teleoperation data.
-“Not Feasible” tasks denote cases where teleoperation failed due to extreme dynamics, precision, or reactivity demands.
-Our method consistently achieves higher success rates and shorter completion times.
-Task Method Success Rate Time (s)
-Hammer Nailing DP 0/13 -
-Ours 13/13 11.0
-Meatball Scooping DP 5/12 42.0
-Ours 10/12 12.4
-Pan Flipping - Egg DP Not Feasible -
-Ours 12/12 1.5
-Pan Flipping - Burger Bun DP Not Feasible -
-Ours 9/12 1.9
-Pan Flipping - Meat Patty DP Not Feasible -
-Ours 10/12 2.3
-Wine Balancing DP Not Feasible -
-Ours 8/10 30.9
-Soccer Ball Kicking DP Not Feasible -
-Ours 6/10 2.0
-TABLE III: Task success rates comparing our method with the hand-held gripper-based method on Nail Hammering.
-Method Demo Duration & Count Success Rate
-UMI [16] $\sim$180 seconds (25 demos) 0/13
-UMI $\sim$720 seconds (100 demos) 13/13
-Ours $\sim$180 seconds (40 demos) 13/13
-VI Experiment Results
-In our experiments, we demonstrate that our framework is both effective and efficient for training robots with advanced capabilities.
-Furthermore, leveraging human play data enables robots to perform smoother movements and acquire skills that are challenging or even impossible to achieve with robot-generated data.
-Figure 5: Policy Testing Across Camera Poses in Nail Hammering.
-(a) Camera poses for data collection and evaluation.
-(b-d) Performance ranges for methods trained with/without random cropping (RC) and view augmentation (VA).
+Figure 10: Fast Tool Changer. Two designs are shown: the left accommodates general tools with a screw mechanism, and the right clips onto tools with specific mounting shapes.
+Figure 11: Initial States for All Evaluation Episodes.
+All methods are evaluated using the same set of manually defined initial states, overlaid in the image.
+These states ensure diverse variations to test the policy’s spatial generalization capabilities.
+B.2 Implementation Details
+B.2.1 Hardware Design
+We designed two fast tool changers compatible with robots using the ISO 9409-1-50-4-M6 flange, as shown in Figure 10.
+The left design utilizes a screw mechanism to accommodate general tools, while the right design employs clips for tools with specific mounting shapes.
+B.2.2 Tool Pose Estimation
+We use Polycam to scan the tool and obtain its mesh. The mesh is later feed into FoundationPose [62] for 6D pose estimation.
 Figure 12: Policy Execution Trajectory Comparison.
@@
 (b) Comparison of end-effector XY trajectories from our framework and a policy trained on robot-collected data.
-Our method generates smoother, more natural, and efficient trajectories with fewer waypoints, enabling fluid and precise motion.
-In contrast, the robot-collected policy produces jerky movements with excessive waypoints, which hinder task performance.
-VI-A Capabilities and Effectiveness
-Table II presents the results of our real-world robot tasks, showing that our framework consistently outperforms baseline methods by achieving significantly higher success rates across all evaluated scenarios.
-We further compare our method with a stronger hand-held gripper baseline, UMI [16], as shown in Table III.
-In our default setup, SLAM-based mapping failed due to low environmental texture.
-To address this limitation, we added a textured background to support reliable mapping for UMI.
-For the nail hammering task, we evaluated UMI using 25 demonstrations (matching our collection time) and 100 demonstrations (to assess its ideal performance).
-UMI failed all 13 trials with 25 demonstrations but succeeded in all 13 trials with 100.
-UMI could not be applied to the wine balancing task due to contact-induced tool displacement, nor to the pan flipping task due to tool inertial slippage.
-In the soccer kicking task, large and fast motions made it nearly impossible to localize the demonstration trajectory within the initial map.
-In contrast to the baselines, our method demonstrates reliable performance across all tasks.
-As illustrated in Figure 4, our method excels in real-world policy rollouts by accurately detecting spatial locations in tasks such as nail hammering and meatball scooping.
-For pan flipping tasks, it performs the high-speed motions required to effectively flip eggs, burger buns, and meat patties.
-Additionally, it demonstrates precise control in lifting and inserting a wine bottle into its stand and reacts swiftly and accurately to kick a soccer ball in the soccer ball kicking task.
-The superior performance of our framework is primarily due to its ability to collect a significantly larger volume of episodes within the same data collection timeframe, as detailed later in Section VII.
-This broader dataset covers a wider range of task variations, enabling more robust and adaptable policy training.
-Furthermore, our approach overcomes the inherent limitations of Gello and SpaceMouse, enabling the collection of demonstrations for scenarios that these tools cannot adequately handle due to their constraints.
-VI-B Policy Execution Efficiency
-Our framework demonstrates exceptional efficiency in task execution, achieving faster task completion times and producing smoother action motions compared to baseline methods, as shown in Table II.
-The efficiency is largely attributed to the nature of human play data, which captures the fluidity and speed of real-world human activities, resulting in smoother and more natural trajectories in the training dataset.
-In contrast, previous approaches relying on teleoperated data often suffer from significantly slower speeds and less natural motions, limiting their effectiveness in dynamic scenarios.
-By leveraging the realistic dynamics of human play, our framework not only accelerates task execution but also enhances motion quality, making it better suited for real-world applications.
-Figure 7: Initial States for All Evaluation Episodes.
-All methods are evaluated using the same set of manually defined initial states, overlayed in the image.
-These states ensure diverse variations to test the policy’s spatial generalization capabilities.
-Figure 8: Robustness to Camera and Base Movement.
-(a) Camera Pose Robustness: The policy demonstrated the ability to handle camera shaking across three tasks—meatball scooping, nail hammering, and pan flipping.
-The first row shows the camera view, while the second row provides a scene overview with the shaking motion.
-(b) Robot Base Robustness: The policy successfully compensated for base shaking, even when the shaking frequency exceeded the robot’s control frequency.
-(c) Chicken Head Stabilization: At lower base movement frequencies, the end effector displayed a stabilization effect similar to a chicken’s steady head.
-(d) Combined Robustness: The policy maintained task performance under simultaneous camera and base shaking.
-Figure 9: Human Perturbation Robustness.
-This figure showcases the robot’s ability to handle human-induced perturbations across three tasks: (1) In nail hammering, the robot successfully followed a manually moved nail; (2) In meatball scooping, it located and scooped meatballs even when additional ones were thrown into the pan mid-task; and (3) In egg flipping, the robot consistently flipped the egg back after human intervention repositioned it.
-These results highlight the policy’s robustness to unpredictable human perturbations.
-VI-C Benefits of Tool-Based Action Representation in Task Space
-We observed that using the tool in the camera frame as the action representation allows the policy to perform comparably to prior works when the camera is stationary.
-However, the success rate drops to zero when the camera is moving.
-Tracking the camera pose in real time during motion is non-trivial, leading to incorrect calculations of the end-effector position in the base frame.
-When the base moves, policies relying on the tool in the base frame fail entirely, achieving zero success.
-These failures occur because such policies assume a fixed relative transformation between the base and the workspace, an assumption invalidated by base motion.
-In contrast, representing actions using the tool in task space enables the policy to remain effective under base perturbations.
-Our approach demonstrates robust performance when camera poses vary, when the base moves, and even during simultaneous shaking of the camera and base.
-VI-D Effects of Random Cropping and View Augmentation
-Our experiments show that random cropping (RC) and view augmentation (VA) enhance policy robustness to diverse camera poses.
-Training with RC improves resilience to minor camera perturbations, such as small movements or shaking, while VA extends robustness by exposing the model to a broader range of camera configurations during training.
-We evaluated these techniques on the nail hammering task and present the results in Figure 5, comparing three models: one trained with both RC and VA, one with RC only, and one without either technique.
-The combined use of RC and VA significantly expands the working range of camera configurations, including those with substantial deviations from the original viewpoint.
-The robustness arises from the model’s exposure to a diverse set of camera poses during training.
-VI-E Policy Execution Trajectory Comparison
-Our framework generates faster, smoother, and more natural trajectories compared to traditional approaches, as demonstrated by the end effector (EEF) XY trajectory for the meatball scooping task in Figure 6.
-Figure 6(a) depicts the initial setup of one trial, while Figure 6(b) contrasts the XY trajectory of the policy rollout using our framework with that of a policy trained on robot-collected data.
-Our framework surpasses the baseline approach in two critical aspects:
-  • Improved Motion Fluidity: Trajectories generated by our framework are significantly smoother and less jerky, enabling more fluid and precise task execution.
-In contrast, the robot-collected data results in jerky motions caused by excessive waypoints, which disrupt movement precision and degrade performance.
-  • Efficient and Effective Learning: The trajectories produced by our framework contain 10 times fewer waypoints than those from robot-collected data.
-The reduction simplifies the learning space, mitigates cumulative error growth, and minimizes the risk of distribution shifts commonly observed in behavior cloning.
-As a result, our policies exhibit significantly improved sample efficiency.
-VI-F Generalization
-Spatial Generalization: Our policy demonstrates strong generalization across various spatial configurations.
-Figure 7 illustrates the range of initial states tested for multiple tasks.
-In the hammer nailing task, nail positions vary.
-For meatball scooping, meatball positions are tested.
-In soccer ball kicking, we evaluate two goalkeeper configurations.
-For pan flipping, object poses vary to test all areas of the pan.
-As shown in Table II, our method achieves high success rates across these ranges.
-Object Generalization: Our method generalizes effectively to different objects in pan flipping tasks, including a toy egg, a 3D-printed meat patty, and a real burger bun (Figure 7, second column).
-With only 13 demonstrations, the policy succeeds by leveraging a simple but effective strategy: tilting the pan to slide the object into a corner, then flicking the pan to propel and flip the object.
-Our manipulation approach for pan flipping enables robust generalization across diverse object types.
-Tool Generalization: To assess the generalization ability of our policy across different tools, we conducted a pan-flipping experiment using a burger bun and five pans: large, medium, small, tiny, and square.
-For each pan, we collected 12 trials with varying initial configurations and reported the success rates (Fig.
-10).
-The policy was trained on demonstrations using the large, medium, and square pans, and evaluated on all five.
-Results indicate that our method exhibits some generalization across both pan sizes and shapes (circular vs.
-square).
-High success rates were observed with the large and medium pans.
-However, performance declined on smaller pans, likely due to their limited surface area.
-The square pan also showed lower success rates, as its shallow edges allowed the bun to slide out during flipping.
-Figure 10: Tool Generalization (a) The tested pans. (b) Success rate across 12 testing trials.
-TABLE IV: Comparison of Data Collection Methods.
+Figure 13: Tool Generalization. (a) The tested pans. (b) Success rate across 12 testing trials.
+Figure 14: Effects of Embodiment Segmentation.
+(a) Failure cases without segmentation: In the wine balancing task, the robot strikes the table, triggering safety stops.
+In the soccer ball kicking task, it performs shorter, less precise actions.
+(b) Quantitative results: Segmentation improved success rates in wine balancing (8 vs.
+0) and soccer ball kicking (6 vs.
+2) by reducing the visual gap between training and deployment.
+Appendix C Additional Experimental Results
+[…]
+Effects of Embodiment Segmentation: Embodiment Segmentation masks the agent’s embodiments during data collection and policy deployment, ensuring visually consistent scenes and reducing the training-deployment visual gap.
+Embodiment Segmentation significantly improves policy performance, as shown in Figure 14.
+Figure 14(a) highlights failure cases without segmentation.
+In the wine balancing task, the robot strikes the table, triggering safety stops due to improper bottle handling.
+In the soccer ball kicking task, the robot’s actions are inconsistent, shorter, and less precise than during training.
+Quantitative results in Figure 14(b) further underscore segmentation’s impact.
+Across 10 trials, segmentation enabled 8 successes in the wine balancing task, while the model without it achieved none.
+Similarly, in the soccer ball kicking task, segmentation resulted in 6 successes, compared to 2 without it.
+By aligning training and testing visual distributions, Embodiment Segmentation ensures consistent and reliable robot performance during the training and deployment.
+Appendix D Detailed Analysis on Data Collection Efficiency and Affordability
+We compare various data collection methods for robot imitation learning, focusing on throughput, reliability, cost, usability, and precision.
+Our evaluation includes teleoperation tools like Gello and Spacemouse for 6DOF (UR5e) and 7DOF (Kinova Gen3) robots, alongside methods such as Visual Imitation Made Easy, handheld grippers (e.g., UMI and LEGATO), and devices like VR (Meta Quest 2), AR (Apple Vision Pro), and Kinematic replicate (Gello).
+D.1 Data Collection Efficiency
+Our framework achieves significantly higher data collection throughput than traditional methods, enabling more demonstrations within the same timeframe.
+The improvement is driven by the natural and intuitive efficiency of human manipulation, which ensures faster and more reliable task execution.
+Figure 7(a) highlights the superior manipulation capabilities of human hands, while Figure 8 quantifies the substantial time savings per episode.
+For nail hammering and meatball scooping, Gello and Spacemouse were used as teleoperation methods, respectively.
+Human hands reduced data collection time by 73% and 81% for nail hammering and meatball scooping, with consistently low variation in performance.
+In more complex tasks like pan flipping, wine balancing, and soccer ball kicking, teleoperation methods failed entirely due to limitations such as lack of tactile feedback, delays, and difficulty handling dynamic or precise actions.
+Our method further reduces data collection time by 41% compared to handheld grippers such as UMI [14] in nail hammering.
+UMI proved ineffective in wine balancing and pan flipping due to tool inertial slippage or contact-induced displacement, and failed in soccer kicking because of difficulty localizing large, fast motions.
+Moreover, it requires rich textures to build a pre-collection map, which our method does not.
+These results underscore the superior efficiency, robustness, and versatility of human manipulation as a scalable solution for high-quality robot learning datasets.
+D.2 Reliability
+Figure 7(b) and Figure 7(c) illustrates typical failure cases with Gello, Spacemouse, and UMI [14], which frequently encounter issues such as safety stops or collisions during data collection.
+In contrast, our method ensures smooth, uninterrupted operation, avoiding these limitations.
+Traditional methods face significant challenges in high-speed or complex tasks.
+For example, Gello and Spacemouse struggle with replicating the extreme dynamics and precise motions required for flipping objects like eggs during pan flipping, often resulting in unsuccessful attempts.
+Similarly, teleoperation delays prevent timely strikes during soccer ball kicking, consistently leading to missed kicks and repeated failures.
+In tasks like wine balancing, the absence of tactile feedback impairs precision during the data collection, causing the wine bottle to tip over during data collection.
+Furthermore, in meatball scooping, the velocity vectors generated by Spacemouse input lead to jerky trajectories with redundant waypoints, significantly reducing efficiency.
+These challenges make effective training impractical with traditional methods.
+By leveraging human manipulation, our framework not only addresses these limitations but also provides a reliable and scalable solution for dynamic and precision-demanding tasks.
+D.3 Discussion of Data Collection Methods
+Table 4 compares various data collection methods based on cost, usability, expertise requirements, intuitiveness, and precision.
+Our method incurs no additional cost (\$0), unlike hardware-dependent solutions like UMI and LEGATO, which demand significant investment.
+This affordability makes our approach accessible to users from diverse backgrounds without financial constraints.
+Unlike hardware-based systems such as UMI, LEGATO, Gello, and Spacemouse, which are prone to malfunctions and maintenance issues, our hardware-free framework ensures reliability and eliminates repair delays or expenses.
+Additionally, it requires no supplementary 3D printing, in contrast to approaches like Visual Imitation Made Easy, UMI, and LEGATO.
+The simplicity of our design promotes inclusivity in collecting large-scale dataset for robot learning research.
+Our method also offers a more natural experience compared to tools like Spacemouse, while being far more cost-effective than VR and AR devices.
+Moreover, systems like Gello and Spacemouse lack the precision necessary for dynamic tasks, a limitation addressed by our approach.
+Overall, our method is a cost-effective, and accessible solution for data collection, overcoming key drawbacks of existing approaches while reducing complexity and maintenance needs.
+Table 4: Comparison of Data Collection Methods.
 This table compares various data collection methods for robotics.
@@
 Ours \$0 Yes No Yes Minimal
-Figure 11: Data Collection Efficiency and Reliability.
-(a) Human hands excel in manipulation tasks, leveraging natural and intuitive efficiency.
-(b) Failure cases for Gello and Spacemouse include insufficient speed, lack of tactile feedback during data collection, safety stops, collisions, teleoperation delays, and difficulty handling high-speed or complex tasks.
-(c) Failure cases for handheld grippers such as UMI [16], where issues arise from tool slippage due to inertia or displacement caused by contact forces.
-Figure 12: Quantitative Comparison of Data Collection Methods.
-Human hands significantly reduce data collection time—by 73% for nail hammering and 81% for meatball scooping—while maintaining consistently low variation across trials.
-Teleoperation fails entirely in dynamic tasks and tasks requiring high precision.
-In nail hammering, human hands are 41% faster than UMI [16], which also struggles in dynamic scenarios and low-texture environments.
-VI-G Robustness
-Camera Pose Robustness: We evaluated the policy’s ability to handle camera pose variations by introducing camera shaking in three tasks: meatball scooping, nail hammering, and pan flipping (egg), as shown in Figure 8(a).
-The first row shows the camera view, while the second row shows the scene overview and the shaking motion.
-Despite the disturbances, the policy successfully completed all tasks.
-This robustness stems from random cropping during training, which allows the policy to adapt to partially cropped inputs and minor visual changes.
-Robot Base Robustness: To evaluate robustness to base movement, we manually shook the robot base during task execution (Figure 8(b)).
-When the shaking frequency exceeded the control frequency, the end effector oscillated with the base.
-However, the task-space action design enabled the robot to compensate for these movements and maintain task success.
-Additionally, as shown in Figure 8(d), the policy remained effective even when both camera and base shaking were applied simultaneously.
-Chicken Head Stabilization: At lower frequencies of base movement, where the shaking was slower than the robot’s control frequency, the end effector exhibited a stabilization effect similar to a chicken’s head remaining steady [75], as shown in Figure 8(c).
-The behavior highlights the robot’s ability to maintain precise control during mild perturbations.
-Human Perturbation Robustness: We further tested the policy’s resilience to human-induced perturbations across three tasks, as shown in Figure 9.
-In the nail hammering task, the robot successfully tracked and followed a manually moved nail by a human.
-During meatball scooping, the robot located and scooped meatballs, even when additional meatballs were thrown into the pan mid-task.
-For egg flipping, the robot consistently flipped the egg back each time a human intervened and returned the egg to different initial positions.
-These results underscore the policy’s robustness in maintaining reliable performance under real-time human interactions.
-VII Data Collection Efficiency and Affordability
-We compare various data collection methods for robot imitation learning, focusing on throughput, reliability, cost, usability, and precision.
-Our evaluation includes teleoperation tools like Gello and Spacemouse for 6DOF (UR5e) and 7DOF (Kinova Gen3) robots, alongside methods such as Visual Imitation Made Easy, handheld grippers (e.g., UMI and LEGATO), and devices like VR (Meta Quest 2), AR (Apple Vision Pro), and Kinematic replicate (Gello).
-VII-A Data Collection Efficiency
-Our framework achieves significantly higher data collection throughput than traditional methods, enabling more demonstrations within the same timeframe.
-The improvement is driven by the natural and intuitive efficiency of human play, which ensures faster and more reliable task execution.
-Figure 11(a) highlights the superior manipulation capabilities of human hands, while Figure 12 quantifies the substantial time savings per episode.
-For nail hammering and meatball scooping, Gello and Spacemouse were used as teleoperation methods, respectively.
-Human hands reduced data collection time by 73% and 81% for nail hammering and meatball scooping, with consistently low variation in performance.
-In more complex tasks like pan flipping, wine balancing, and soccer ball kicking, teleoperation methods failed entirely due to limitations such as lack of tactile feedback, delays, and difficulty handling dynamic or precise actions.
-Our method further reduces data collection time by 41% compared to handheld grippers such as UMI [16] in nail hammering.
-UMI proved ineffective in wine balancing and pan flipping due to tool inertial slippage or contact-induced displacement, and failed in soccer kicking because of difficulty localizing large, fast motions.
-Moreover, it requires rich textures to build a pre-collection map, which our method does not.
-These results underscore the superior efficiency, robustness, and versatility of human play as a scalable solution for high-quality robot learning datasets.
-VII-B Reliability
-Figure 11(b) and Figure 11(c) illustrates typical failure cases with Gello, Spacemouse, and UMI [16], which frequently encounter issues such as safety stops or collisions during data collection.
-In contrast, our method ensures smooth, uninterrupted operation, avoiding these limitations.
-Traditional methods face significant challenges in high-speed or complex tasks.
-For example, Gello and Spacemouse struggle with replicating the extreme dynamics and precise motions required for flipping objects like eggs during pan flipping, often resulting in unsuccessful attempts.
-Similarly, teleoperation delays prevent timely strikes during soccer ball kicking, consistently leading to missed kicks and repeated failures.
-In tasks like wine balancing, the absence of tactile feedback impairs precision during the data collection, causing the wine bottle to tip over during data collection.
-Furthermore, in meatball scooping, the velocity vectors generated by Spacemouse input lead to jerky trajectories with redundant waypoints, significantly reducing efficiency.
-These challenges make effective training impractical with traditional methods.
-By leveraging human play, our framework not only addresses these limitations but also provides a reliable and scalable solution for dynamic and precision-demanding tasks.
-VII-C Discussion of Data Collection Methods
-Table IV compares various data collection methods based on cost, usability, expertise requirements, intuitiveness, and precision.
-Our method incurs no additional cost (\$0), unlike hardware-dependent solutions like UMI and LEGATO, which demand significant investment.
-This affordability makes our approach accessible to users from diverse backgrounds without financial constraints.
-Unlike hardware-based systems such as UMI, LEGATO, Gello, and Spacemouse, which are prone to malfunctions and maintenance issues, our hardware-free framework ensures reliability and eliminates repair delays or expenses.
-Additionally, it requires no supplementary 3D printing, in contrast to approaches like Visual Imitation Made Easy, UMI, and LEGATO.
-The simplicity of our design promotes inclusivity in collecting large-scale dataset for robot learning research.
-Our method also offers a more natural experience compared to tools like Spacemouse, while being far more cost-effective than VR and AR devices.
-Moreover, systems like Gello and Spacemouse lack the precision necessary for dynamic tasks, a limitation addressed by our approach.
-Overall, our method is a cost-effective, and accessible solution for data collection, overcoming key drawbacks of existing approaches while reducing complexity and maintenance needs.
-VIII Limitations and Future Work
-Our framework has certain limitations.
-First, the perception pipeline relies on FoundationPose for extracting the tool’s pose during manipulation.
-Errors in pose estimation may occasionally require data recollection, adding time and effort.
-Improving the reliability of the perception pipeline through more robust pose estimation algorithms or self-correction mechanisms is a promising direction for future work.
-Second, for novel view augmentation, significant noise and reduced realism are observed when augmented views deviate too far from the collected camera views, which can hinder policy performance.
-Future efforts could focus on leveraging advanced rendering techniques to enhance the realism of augmented views and improve policy generalization.
-Third, we assume the tool is rigidly attached to the robot’s end effector; however, in real-world, contact-rich manipulation, minor shifts may occur, potentially affecting performance.
-Addressing this issue by incorporating tactile sensing could improve performance in contact-intensive tasks.
-Additionally, our method assumes a rigid tool and does not account for flexible or soft tools.
-Future work could explore using flexible representations for tool state estimation to better handle deformable tools in real-world manipulation scenarios.
-IX Conclusion
-In this work, we presented a novel framework for human-to-robot imitation learning that leverages human play data to bridge the embodiment gap and enables robust policy training for diverse tool-use tasks.
-Unlike traditional data collection methods, which are often costly, hardware-dependent, and require technical expertise, our framework democratizes data collection by removing the need for specialized equipment or prior knowledge.
-Our approach makes data collection more accessible and scalable, empowering broader adoption in robotic learning.
-We validated our framework across a range of challenging tasks, including nail hammering, meatball scooping, pan flipping with various objects, wine bottle balancing, and soccer ball kicking.
-The results demonstrate the framework’s superior performance, robustness to variations in camera poses and base movements, and adaptability to different embodiments, such as 6-DOF and 7-DOF robots.
-By enhancing accessibility, scalability, and reliability, our work lays a strong foundation for advancing robotic manipulation in complex, real-world scenarios.
-Acknowledgments
-We thank Professor Wenzhen Yuan, Xiaoyu Zhang, Yucheng Mo, Yilong Niu, Hyoungju Lim, and Amin Mirzaee for their support and assistance in reproducing the UMI.
-We thank Professor Saurabh Gupta, Professor Junyi Geng, Shuijing Liu, Lujie Yang, Neeloy Chakraborty, Hongkai Dai, Jacob Wagner, Shaoxiong Yao, Wei-Cheng Huang for their insightful feedback and suggestions.
-This work was supported by ZJU-UIUC Joint Research Center Project No.
-DREMES 202003, funded by Zhejiang University.
-This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
-Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
-Additionally, this work used NCSA Delta GPU at NCSA through allocation CIS240753 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S.
-National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
-Appendix A Appendix
-A-A Implementation Details
-A-A1 Hardware Design
-We designed two fast tool changers compatible with robots using the ISO 9409-1-50-4-M6 flange, as shown in Figure 13.
-The left design utilizes a screw mechanism to accommodate general tools, while the right design employs clips for tools with specific mounting shapes.
-Figure 13: Fast Tool Changer. Two designs are shown: the left accommodates general tools with a screw mechanism, and the right clips onto tools with specific mounting shapes.
-A-A2 Tool Pose Estimation
-We use Polycam to scan the tool and obtain its mesh. The mesh is later feed into FoundationPose [71] for 6D pose estimation.
-A-B Additional Results
-A-B1 Effects of Embodiment Segmentation
-Embodiment Segmentation masks the agent’s embodiments during data collection and policy deployment, ensuring visually consistent scenes and reducing the training-deployment visual gap.
-Embodiment Segmentation significantly improves policy performance, as shown in Figure 14.
-Figure 14(a) highlights failure cases without segmentation.
-In the wine balancing task, the robot strikes the table, triggering safety stops due to improper bottle handling.
-In the soccer ball kicking task, the robot’s actions are inconsistent, shorter, and less precise than during training.
-Quantitative results in Figure 14(b) further underscore segmentation’s impact.
-Across 10 trials, segmentation enabled 8 successes in the wine balancing task, while the model without it achieved none.
-Similarly, in the soccer ball kicking task, segmentation resulted in 6 successes, compared to 2 without it.
-By aligning training and testing visual distributions, Embodiment Segmentation ensures consistent and reliable robot performance during the training and deployment.
-Figure 14: Effects of Embodiment Segmentation.
-(a) Failure cases without segmentation: In the wine balancing task, the robot strikes the table, triggering safety stops.
-In the soccer ball kicking task, it performs shorter, less precise actions.
-(b) Quantitative results: Segmentation improved success rates in wine balancing (8 vs.
-0) and soccer ball kicking (6 vs.
-2) by reducing the visual gap between training and deployment.